### v6

# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

Abstract
Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily. Efficiently managing, processing, and displaying this data requires specialized expertise and is often tedious and repetitive. Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution. However, LLMs are not adept at handling complex numerical computations and table manipulations and are also constrained by a limited context budget. Based on this, we propose Data-Copilot, a data analysis agent that autonomously performs querying, processing, and visualization of massive data tailored to diverse human requests. The advancements are twofold: First, it is a code-centric agent that receives human requests and generates code as an intermediary to handle massive data, which is quite flexible for large-scale data processing tasks. Second, Data-Copilot involves a data exploration phase in advance, which explores how to design more universal and error-free interfaces for real-time response. Specifically, it actively explores data sources, discovers numerous common requests, and abstracts them into many universal interfaces (code modules) for daily invocation. When deployed in real-time requests, Data-Copilot only needs to invoke these pre-designed interfaces, transforming raw data into visualized outputs (e.g., charts, tables, and text) that best match the user’s intent. Compared to generating code from scratch, invoking these pre-designed and compiler-validated interfaces can significantly reduce errors during real-time requests. Additionally, interface workflows are more efficient and offer greater interpretability than code. We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news, demonstrating promising application prospects. Quantitative evaluations indicate that, compared to direct code generation and other agent strategies, our exploration-deployment strategy addresses human requests more accurately and efficiently, i.e., with higher success rates and lower token consumption.
Figure 1: We present two strategies for using LLMs to automate data analysis. Upper: Directly using LLMs to read and process massive data is limited by context length, the LLM’s capabilities, and output format constraints. Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks. It actively explores how to design more universal and error-free interface modules, improving the success rate of real-time requests. Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow to fulfill human intent.
1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption. However, these data-related tasks often require specialized knowledge and tedious manual labor.
Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022), especially when combined with techniques such (Wei et al., 2022b; Kojima et al., 2022) has demonstrated their capability to handle complex tasks. Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow that performs data analysis and outputs visualization results that best match user expectations?
An intuitive solution is to directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023), just as we do with text. However, several challenges must be considered: (1) Due to the input context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text. Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources. (2) Data processing is very complex, involving many tedious numerical calculations and intricate table manipulations. LLMs are not adept at performing these tasks. (3) An LLM can only output text, whereas we usually prefer charts or tables when analyzing data. These challenges significantly constrain the application of LLMs in data-related tasks.
Recently, many works have explored alternative solutions. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024) and DB-GPT (Xue et al., 2023) apply LLMs to data science domain like Text2SQL. Besides, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). Data Interpreter (Hong et al., 2024) proposes a code-based agent for data science through the plan-code-verify paradigm. These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent. Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and output multi-modal results (e.g., chart, table, text) for humans, reducing tedious human labor and eliminating the need for direct data access by the LLM. However, creating a reliable code agent for various data analysis tasks is also not an easy feat. Firstly, LLM-generated code often contains formatting errors, logical inconsistencies, and code hallucinations, such as fabricating non-existent functions, which may prevent it from being deployed in real-world environments. Secondly, most code-based agent generates code for solving user requests, without considering the relations between different requests and only addressing each issue independently.
Based on this observation, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis. First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches the user’s intent. The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks. Besides, Data-Copilot also incorporates a data exploration phase. During this phase, it actively explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance. In a real-world deployment, Data-Copilot flexibly invokes pre-design interface modules to handle user requests, creating an interface workflow that transforms raw data into desired visualization results.
This design can bring three advantages. Firstly, this exploratory process allows Data-Copilot to analyze connections between data, design general interfaces for similar requests, and pre-validate their correctness, thereby reducing errors in real-time responses. Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving generation efficiency. Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
To achieve this, as shown in Figure 1, Data-Copilot contains three steps for data exploration:
  1. 1.
    Request Exploration: Data-Copilot explores potential requests and mines a broader range of human need scenarios from data. It involves a "self-instruct" process to generate massive simulated requests based on all data schemas (types and formats) and seed requests.
  2. 2.
    Interface Design: Our agent generates code for each request and then considers whether these requests have similar demands and whether general interfaces can be abstracted.
  3. 3.
    Interface Optimization: To improve the quality of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces by reflecting on feedback from the compiler.
After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different request scenarios, Data-Copilot can flexibly design different interface invocation procedures, such as serial workflows, i.e., step-by-step invocation, parallel workflows, and loop workflows. Our contributions are threefold:
Figure 2: Overview of Data-Copilot. Data Exploration: It devises a self-request process to generate numerous potential requests based on data schemas. Then it interleaves to design and optimize the interfaces to solve these requests. The whole exploration-design-optimization process is operated in advance, generating many general and error-free interface modules, such as data retrieval, computation, and visualization. Workflow Deployment: Upon receiving real-time requests, LLM invokes these interfaces, forming an interface workflow with different structures for request solving.
  1. 1.
    We propose a code-centric agent, Data-Copilot, for massive data analysis tasks. It leverages LLM and its code generation capabilities to automate data querying, processing, and visualization tailored to human requests, reducing tedious human labor.
  2. 2.
    We decouple code generation into two phases: data exploration and workflow deployment. In the exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data. These interfaces enhance the success rate and efficiency of real-time responses. In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests.
  3. 3.
    We open-source Data-Copilot for diverse Chinese financial data analysis, including stocks, funds, and live news. Quantitative evaluations indicate that our agent outperforms other agent strategies, with higher success rates and lower token consumption. Besides, interface workflows are more convenient for human inspection and interaction, offering stronger interpretability.
2 Data-Copilot
Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent. Data-Copilot operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-request process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these "self-instruct" requests, it generates code and abstracts universal interface modules for each request (Step 2: design). After that, it optimizes similar interfaces and reflects on the code compiler feedback, ensuring each interface is universal and correct (Step 3: optimization). Finally, Data-Copilot acquires many generic, error-free interfaces for subsequent use. The whole exploration-design-optimization process is operated in advance.
When faced with real-time requests (Section 2.2), Data-Copilot adeptly comprehends user intentions, and invokes pre-designed interfaces for request solving, significantly reducing the need for tedious labor. We provide a detailed prompt for two phases in Section F.1 and F.2.
2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe what data is available, understand the data formats, and determine how to use them. Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability. Similar to humans, Data-Copilot aims to explore data and derive insights from vast data sources. These insights include understanding the relations between the data, discovering the potential requests associated with the data, and organizing the code framework efficiently. Based on this, Data-Copilot summarizes these exploratory insights into many reusable code components and optimizes its performance. It significantly reduces the code complexity for real-time responses, thus improving the success rate.
Request Exploration
To explore data and mine insights, we design a self-request process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs must generate requests specifically related to these data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.
To achieve this, we first generate a parsing file for each data source to help LLM understand the data. Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example. Then we feed the parsing files and a few seed requests into the LLMs, and prompt LLMs to generate more diverse requests based on these data, covering different demand scenarios as comprehensively as possible. The brief prompt is shown as follows:
###Instruction: Given some data and its description, please mimic these seed requests and generate more requests. The requests you generate should be as diverse as possible, covering more data types and common needs.
###Seed Request: {request1, request2,...}
###Parsing file for GDP_Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: pro.cn-gdp(start-time, end-time, frequency,...),
Output Schema: Return 9 columns, including quarter, gdp, gdp-yoy...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’,..),
First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
###Parsing file for Stock_data:{...},
How to ensure the quality of the generated requests? A common issue is that the LLM often proposes a request about the data that doesn’t exist. To address this, we design a backward verification process to check the rationality of these requests. Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests. Upon manually check some requests, we find that the generated requests generally met our expectations.
Interface Design
After generating massive requests through the self-request, Data-Copilot considers abstracting universal interfaces from these requests. First, we need to define what an interface is in our paper. Similar to human programming practices, we refer to a function module abstracted from the code as an interface. Each interface is composed of a name, parameters, a functional description, and the implementation code, performing specific functions such as data retrieval, computing, and plotting. To abstract general interfaces from all requests, we design an interwoven process, performing interface design and interface optimization alternately for each request.
Firstly, we feed related data parsing files and existing interfaces stored in the interface library (empty at first iteration) into LLMs as a prompt. LLMs need to generate a complete code for each request. When generating, LLMs are instructed to prioritize using the existing interfaces from the previous iterations. If these interfaces are insufficient to address the request, the LLMs design a new interface. As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...). As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
Interface Optimization
To improve the quality of interfaces, Data-Copilot optimizes them in two ways: merging similar interfaces to make them more universal and reflecting on compiler feedback to reduce errors. This optimization process is intertwined with the interface design for each request.
Firstly, similar to software developers wrapping similar code modules to reduce redundancy, Data-Copilot considers whether each newly designed interface can be merged with the previous interfaces. LLMs are instructed to check the functionality, parameters, and logic of the new and existing interfaces, and merge those similar interface modules to create a new one. As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...). This interface optimization process makes each interface more general and rational.
Besides, to reduce code hallucinations or formatting issues, Data-Copilot also self-reflects on the feedback of code execution for each request if compilation errors occur. We have LLMs perform self-reflection until the code passes or the maximum retry number is reached. This process ensures the correctness and usability of each interface.
For each request, Data-Copilot alternates the above two steps: interface design and optimization until all the requests can be satisfied by these interfaces. Finally, as shown in Figure G3, Data-Copilot produces five types of interfaces: data acquisition, processing, prediction, visualization, and DataFrame manipulation. We provide some detailed cases in Section G.1. Besides, to handle newly added data sources, Data-Copilot can repeat the aforementioned exploration-design-optimization process on new data. It can create some new interface modules for these new data.
2.2 Workflow Deployment
In the first phase, Data-Copilot designs many generic interfaces with different functionalities. They are versatile and easy to use, simplifying the code generation of real-time responses. As shown in Figure 2, when faced with human requests, Data-Copilot only needs to invoke related interfaces and form an interface workflow. A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex call flow diagram. This workflow transforms raw data into visualized results. The detailed prompts and cases are shown in Sections F.2 and G.2.
Intent Analysis
To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks. For example, the request is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2019 to May 2024 for comparison". To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
Planning Workflow
After providing the user request and a description of the relevant interfaces, Data-Copilot plans a logical interface workflow to process the user request. Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot. The LLMs are instructed to output in JSON format for each step invocation: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}.
For different requests, Data-Copilot designs interface workflows with different structures, such as sequential, parallel, or loop. For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot plans a five-step workflow. In the first two steps, it dispatches Data Acquisition(), Calcuate Return() in parallel for three indices. The last three steps successively invoke the Visualization() to sequentially plot three indices. In the second case, it deploys a loop workflow by calling the Loop().
Multi-form Output
Upon execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text. Additionally, it also provides a comprehensive summary of the entire workflow. It greatly enhances the interpretability of the whole process, as the interface workflow is easy for humans to read and inspect. As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...". Data-Copilot first interprets the user’s intent. Then it deploys a three-step workflow: 1) Invoking get-GDP-data() interface to acquire historical GDP data. 2) Invoking predict-next-value() interface for forecasting. 3) Visualizing the output.
Table 1: Left: Statistics on human-proposed requests and self-request dataset. Right: Detail Statistics on different request types and workflow structures.
Source #Cases Form Usage
Human-proposed 173 Query Seed
Self-Request 3000 Query Interface Design
Self-Request + Human-annotated 547 Query, Ans Testing
Request Types Workflow Types
Sequential Parallel Complex Overall
Stock 79 106 87 272
Fund 55 36 22 113
Corp. 97 30 4 131
Other 6 12 13 31
Total 237 184 126 547
3 Qualitative Evaluation
3.1 Environment and Data Sources
We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies. The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance. In the first phase, we use the gpt-4-0613 for data exploration design and gpt-3.5-turbo-0613 for workflow deployment.
3.2 Seeds and Datasets
Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each. These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate. 2) involves different levels of difficulty. Then we filter out highly similar requests by computing the sentence embedding of each request. Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
en Self-request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request to expand the request set. Specifically, we feed 173 human-proposed requests along with all data descriptions and schema into GPT-4. We instruct GPT-4 to generate massive requests. In this phase, we generate 6480 requests, then compute the semantic similarity between all requests, and filter out those with high similarity, and retain 3547 requests. We randomly select 3000 requests from the total set for subsequent interface design. Meanwhile, the remaining 547 requests are treated as a test set. We manually annotate the final answers, task types, and workflow types for each testing instance. The statistics of three datasets are shown in Tables 1 and B1.
3.3 Qualitative Results
The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request. The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces. For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
Data-Copilot designs versatile interfaces through data exploration. In Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, manipulation, and visualization. These diverse interfaces endow the LLM with the capability to address complex requests. For instance, get-stock-prices-data() interface integrates the functions of acquiring data by day, week, and month, controlled by argument: freq, which is similar to the process of programmer development.
Data-Copilot deploys workflows with different structures. Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests. As shown in Figure 2, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface. It iteratively calculates the financial indicators for each stock. Please refer to Appendix E for more visualization cases.
4 Quantitative Evaluation
We quantitatively assess Data-Copilot in processing real-time user requests, where it could invoke the interfaces designed in the first phase.
4.1 Experiments Settings
Baselines 1. Vanilla LLM (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests. 2. ReAct (Yao et al., 2022): The LLM resolves the user request using neural language and code step-by-step. 3. Reflexion (Shinn et al., 2023): LLM can also reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 4. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. All methods are evaluated on the 547 testing instances (Section 3.2) three times. The detailed prompts of all baselines are shown in Section F.3.
Metrics We design three metrics for solution code and answer: Exec@1 Pass@1 and Efficiency. Exec@1 evaluates whether the code can be executed by a compiler without errors. Pass@1 measures the correctness of the final answer. Considering the diverse output formats, e.g., images, text, numbers, and tables, we employ GPT-4 to automatically assess the answer. If the output is an image, we use gpt-4-vision-preview to compare the generated image to the label. Lastly, we calculate the total number of tokens consumed by the input prompt and output response, which denotes the solving efficiency of each method.
Table 2: We demonstrate the accuracy and efficiency of all methods with gpt-3.5.
Methods Exec@1(%) Pass@1(%) #Token
[…]
Table 3: All samples are divided into three subsets based on its complexity. We compute Pass@1.
Methods Sequential Parallel Complex Overall
Vanilla 42.4% 29.0% 1.6% 28.5%
ReAct 56.3% 48.6% 14.6% 44.1%
Reflexion 67.8% 55.1% 48.1% 59.0%
Multi-agent 63.2% 64.7% 35.5% 57.4%
Data-Copilot 71.8% 70.0% 67.1% 70.2%
4.2 Comparison Results
As shown in Table 3, Data-Copilot surpasses all baselines in success rate (Pass@1), Exec@1, and Efficiency, achieving 11.2% improvements.
Data-Copilot significantly reduces the risk of failure. Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%. In most cases, it merely needs to invoke pre-designed interface modules and generate corresponding arguments to address user requests, notably reducing the risk of errors. In contrast, baseline strategies must construct all code from scratch, making it easy to overlook certain details or make mistakes. Besides, the difference between Pass@1 and Exec@1 indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
Data-Copilot reduces repetitive overhead for common requests. The consumed tokens represent computational overhead. Data-Copilot only consumes 561 tokens, just 30% of the baseline’s consumption. Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments. The primary token consumption lies in the input prompt which contains related interfaces descriptions. In contrast, baseline strategies may consume a lot of tokens since they need to output a complete solving code for each request. In the real world, numerous requests are similar or even repetitive. Data-Copilot can reduce repetitive overhead by abstracting common demands into interfaces. However, baseline methods have to repetitively generate code for each request.
4.3 Analysis
Data-Copilot exhibits a pronounced advantage on more challenging tasks. To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex. Complex workflows mean some intricate cases, mainly involving loops or nested workflows. As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging parallel or complex tasks, our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing. In contrast, our methods can effectively handle these tasks by invoking versatile interfaces and deploying workflow. These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
We can deploy Data-Copilot locally using small LLMs with commendable performance. We analyze the performance of Data-Copilot with different LLMs. In Figure 4, we test GPT family and various open-source models. We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results. For instance, codellama-13B achieves a 50.8% Pass@1, almost similar to the best baseline in Table 3 (Reflexion with gpt-3.5-turbo). This indicates that Data-Copilot can be deployed locally at a relatively low cost while maintaining commendable performance.
4.4 Human Evaluation
To assess whether the solutions align with human preferences, we perform human evaluations across two dimensions: rationality and conciseness. Specifically, we collect the solutions of Data-Copilot and Reflexion. Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix B). We present results in Figure 4. It demonstrates that our solution significantly outperforms the competitor in terms of rationality. Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%. Besides, as for conciseness, Data-Copilot lags behind its competitor, with 27.8% wins against 30% losses. This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are universal, accommodating a wide array of functions, which makes our system less concise. The results of human evaluation are consistent with the previous assessment. Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
4.5 Performance of other programming languages
Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces. We tested three programming languages: Python, C++, and Matlab. The results indicate that Python performed the best (Pass@1: 70.2), followed by C++ (54.2), with Matlab (36.5) yielding the poorest results. Upon examination, we found that Matlab code often suffers from formatting errors or dimensional discrepancies in the data, rendering the program non-executable. We speculate this may be related to a lack of sufficient Matlab code in the pre-training corpus. So ultimately, we opted for Python for Data-Copilot.
Figure 3: Results of Data-Copilot using various LLMs as based model.
LLM Pass@1 (%) #Token
gpt-4 75.6 561.2
gpt-3.5-turbo 70.2 559.3
Llama-2-70B-chat 49.3 623.6
codellama-13B 50.8 507.4
text-davinci-003 47.5 495.9
Llama-2-13B-chat 36.9 443.8
Vicuna-13B-v1.5 33.2 502.8
Figure 4: Win rate for Data-Copilot Vs. Reflexion on human preferences.
5 Conclusion
We propose a code-centric agent, Data-Copilot, to address a wide array of data-related tasks. Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests. It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving. Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
Appendix
Appendix A Experiments Details
During the data exploration phase, we instruct LLM to generate Python code for each interface. Then we abstract interfaces from these requests. To merge the similar interfaces for each iteration, we adopt the following strategy:
1. Firstly, we calculate the semantic similarity between the names of all existing interfaces and the newly designed interface, then sort them by similarity from highest to lowest.
2. Then, we feed the top-10 existing interfaces into the LLM, and instruct the LLM to sequentially check whether they are functionally similar to the new interface, or can be merged using new parameters.
3. Lastly, we output the merged interface. If the LLM deems no merging necessary, the new interface is retained.
Besides, for the workflow deployment phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface. We set the temperature to 0.2 for more consistent results.
Appendix B Human Evaluation Details
We collect the solutions of Data-Copilot and Reflexion on 173 human-proposed requests. Subsequently, we involve experts to evaluate which of the two solving codes is more rational and concise. The criteria for expert are as follows:
Rationality:
1) Accurately identifies data of interest to the user.
2) Processes the data logically.
3) Output data in a format that meets human needs, with aesthetic and concise presentation.
Conciseness:
1) Only extracts and computes data relevant to the users interest.
2) Code is clear, concise, and well-structured.
3) The solution process is devoid of repetition and redundancy, ensuring a straightforward and intuitive approach.
Experts are required to evaluate which of the two solutions is better based on these principles, and then we calculate the win rate of Data-Copilot against the Reflexion.
Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
Appendix C Limitations
Data-Copilot proposes a new paradigm for addressing the data-related task, through LLM. But we want to highlight that it still remains some limitations or improvement spaces:
1) Online Design Interface. The essence of Data-Copilot lies in effective interface design, a process that directly affects the effectiveness of subsequent interface deployments. Currently, this interface design process is conducted offline. Therefore, it is crucial to explore how to design the interface online and deploy it simultaneously. It greatly broadens the application scenarios of Data-Copilot.
2) System stability The interface deployment process can occasionally be unstable. The main source of this instability is because LLM is not fully controllable. Despite their proficiency in generating the text, LLMs occasionally fail to follow the instructions or provide incorrect answers, thus causing anomalies in the interface workflow. Consequently, finding methods to minimize these uncertainties during the interface dispatch process should be a key consideration in the future.
3) Data-Copilot possesses the potential to handle data from other domains effectively. Currently, our focus is on developing an LLM-based agent for the data domain. Due to the limited access to data, we chose the China Financial Data, which includes stocks, futures, finance, macroeconomics, and financial news. Although these data all belong to the financial domain, the data volume is extremely large, and the data schemas are highly different. The corresponding user’s requests are also diverse, which poses a great challenge to the current LLM. However, Data-Copilot has adeptly accomplished this task. Therefore, we believe Data-Copilot also possesses the potential to effectively handle data from other domains.
Appendix D Related Works
In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa [Brown et al., 2020, Chowdhery et al., 2022, Zhang et al., 2022, Zeng et al., 2023, Touvron et al., 2023, Ouyang et al., 2022, OpenAI, 2023, Wei et al., 2022a] have revolutionized the field of natural language processing (NLP). These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning. The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism [Ouyang et al., 2022, Wang et al., 2022b, a].
LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs). Tool-enhanced studies  [Schick et al., 2023, Gao et al., 2022, Qin et al., 2023a, Hao et al., 2023, Qin et al., 2023b] integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools. Several researchers have extended the scope of LLMs to include the other modality [Wu et al., 2023a, Surís et al., 2023, Shen et al., 2023, Liang et al., 2023b, Huang et al., 2023]. In addition, there are many LLM-based agent applications [Xie et al., 2023], such as CAMEL [Li et al., 2023a], AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^3] https://github.com/reworkd/AgentGPT, BabyAGI [^4] https://github.com/yoheinakajima/babyagi, BMTools [^5] https://github.com/OpenBMB/BMTools, LangChain [^6] https://github.com/hwchase17/langchain, Agentverse [Chen et al., 2023b], Autoagent [Chen et al., 2023a], MetaGPT [Hong et al., 2023], AutoGEN [Wu et al., 2023b], etc. Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks. Except for learning to operate the tools, several contemporaneous studies [Cai et al., 2023, Qian et al., 2023] have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning. These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers. FLAME [Joshi et al., 2023] investigates the feasibility of using NLP methods to manipulate Excel sheets. StructGPT [Jiang et al., 2023] explore reasoning abilities of LLM over structured data. LiDA [Dibia, 2023] and GPT4-Analyst [Cheng et al., 2023] focus on automated data exploration. Besides, many reseraches [Liu et al., 2023, Chang and Fosler-Lussier, 2023, Dong et al., 2023], like Sheet-Copilot [Li et al., 2023b], BIRD [Li et al., 2023c], DAIL-SQL [Gao et al., 2023], DIN-SQL [Pourreza and Rafiei, 2023], PET-SQL [Li et al., 2024b], DB-Copilot [Wang et al., 2023b], MAC-SQL [Wang et al., 2023a], ACT-SQL [Zhang et al., 2023], ChatBI [Lian et al., 2024], CodeS [Li et al., 2024a], SQLPrompt [Sun et al., 2023a], ChatDB [Hu et al., 2023], SQL-PaLM [Sun et al., 2023b], and DB-GPT [Xue et al., 2023], EHRAgent [Shi et al., 2024] apply LLMs to Text2SQL and table rasoning. Chain-of-Table [Wang et al., 2024] proposes a step-by-step reasoning strategy based on the table.
Distinct from these approaches, our Data-Copilot system is different in the following aspects:
(1) Different Task Settings: Unlike Text2SQL methods, Data-Copilot focuses more on how to utilize data and present it according to user preferences, rather than retrieving data from databases. We assume that Data-Copilot has access to a vast amount of data, which is available via API.
(2) Richer Usage Scenarios: Data-Copilot needs to meet a diverse range of human needs, such as forecasting and plotting. Therefore, it employs programmable language like Python for data processing, offering more varied use cases than SQL.
(3) Distinct Implementation Approach: Data-Copilot is a code-centric agent for data processing, rather than relying on SQL or pre-existing tools. This approach is more flexible and can meet complex user requirements.
Appendix E Visualization
We provide several cases in this section to visualize workflow deployed by Data-Copilot, which includes queries about diverse sources (stocks, company finance, funds, etc.) using different structures (parallel, serial, and loop Structure).
Different Structures
As shown in Figure G4,G5, Data-Copilot deploys different structural workflows based on user requirements. In Figure G4, the user proposes a complex request, and Data-Copilot deploys a loop structure to implement the financial data query of each stock, and finally outputs a graph and table in parallel. In Figure G5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas. These concise workflows can cope with such complex requests well, which suggests that the data exploration and workflow deployment process of Data-Copilot are rational and effective.
Diverse Sources
Figure G6, G7, G8 demonstrate that Data-Copilot is capable of handling a large number of data sources, including stocks, funds, news, financial data, etc. Although the formats and access methods of these data types are quite different, our system efficiently manages and displays the data through its self-designed versatile interface, requiring minimal human intervention.
Appendix F Detailed Prompts
We provide detailed prompts for our Data-Copilot and baselines. Specifically, for the data exploration, we provide detailed prompts in Section F.1 for four phases: self-request, interface design, and interface optimization. For the workflow deployment phase, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow. Additionally, we outline prompts for all baselines in Section F.3: Vanilla LLM, ReAct, Reflection, and Multi-Agent Collaboration strategies.
F.1 Prompts For Data Exploration
Explore Data by Self-Request Phase
###Instruction: Given some data and its description, please mimic these seed requests and generate more requests. The requests you generate should be as diverse as possible, covering more data types and common needs.
###Seed Request: {request1, request2,...}
###parsing file for GDP-Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: You can access the data by pro.cn-gdp(start-time, end-time, frequency,...), start-time means....,
Output Schema: The data return 9 columns, including quarter: quarter, gdp: cumulative GDP, gdp-yoy: quarterly Year-on-Year growth rate, pi,...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’, frequency=’quarter’),
First Row: {2019Q4 990865.1 6.10, ...},
Last Rows: {2018Q4 900309.5 6.60, ...}}}
###parsing file for stock_data: {...}
Interface Design:
###Instruction: You are an experienced program coder. Given a request and some existing interfaces, you should use these interfaces to solve the request or design new interfaces to resolve my request.
(1) You should define the name, function, inputs, and outputs of the interface. Please describe the functionality of the interface as accurately as possible and write complete implementation code in the new interface.
(2) Finally please explain how to resolve my request using your newly designed interfaces or existing interfaces in the neural language.
###Output Format:
Your newly designed interfaces:
Interface1={Interface Name: {name},
Function description: {This interface is to ...},
Input: {argument1: type, argument2: type, ...},
Output: {pd.DataFrame} }
Interface2=....
The solving process for request: {To fullfil this request, I design a interface ...}
###The user request: {input request}
###Data Files: {All data parsing files}
###The existing interfaces: {all interfaces in library}
Interface Optimization
###Instruction: Please check that the interface you have designed can be merged with any existing interfaces in the library.
(1) You should merge interfaces with similar functionality and similar input and output formats into a new interface.
(2) You can use parameters to control the different inputs, if you want to merge two interfaces.
(3) Please explain your reason for merging and output all interfaces in the library after merging.
(4) If you dont think a merge is necessary, then just add new interfaces into the existing interface library and output them all.
###New interfaces: {interfaces}
###Existing interfaces: {interface1, interface2, interface3, ..}
###Output Format:
The reasons for merging: {reason}
Interfaces after merging: {interface1, interface2, optimized interface3, ..}
F.2 Prompts For Workflow Deployment
Intent Analysis Phase
### Analysis Prompt: Please parse the input request for time, place, object, and output format. You should rewrite the instruction according to todays date. The rewritten new instruction must be semantically consistent and contain a specific time and specific indicators.
### Output Format: Rewritten Request. (Time:%s, Location:%s, Object:%s, Format:%s).
### User Request: Today is {Timestamp}. The user request is {Input Request}.
Please output a Rewritten Request.
Task Selection
###Select Prompt: Please select the most suitable task according to the given Request and generate its task_instruction in the format of task={task_name: task_instruction}. There are four types of optional tasks. [fund_task]: used to extract and process tasks about all public funds. [stock_task]: for extracting and processing tasks about all stock prices, index information, company financials, etc., [economic_task]: for extracting and processing tasks about all Chinese macroeconomic and monetary policies, as well as querying companies and northbound funds, [visualization_task]: for drawing one or more K-line charts, trend charts, or outputting statistical results.
###Output Format: task1={%s: %s}, task2={%s: %s}
###User Request: {Rewritten Request}.
Please output a task plan for this request.
Planning Workflow
###Planning prompt: Please use the given interface (function) to complete the Instruction step by step. At each step you can only choose one or more interfaces from the following interface library without dependencies, and generate the corresponding arguments for the interface, the arguments format should be strictly in accordance with the interface description. The interface in the later steps can use results generated by previous interfaces.
###Output Format:
Please generate as json format for each step:step1={"arg1": [arg1,arg2...], "function1": "%s", "output1": "%s", "description1": "%s"}, step2={"arg1": [arg1,arg2..], "function1": "%s", "output1": "%s", "description1": "%s"}, ending with ###.
###User Request: {Task Instruction}.
Please output an interface invocation for this instruction.
F.3 Prompts For Baselines
Vanilla LLM
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, you should write a complete Python code to fulfill the users request. Your code must completely fulfill all the users requirements without syntax errors!
###User Request: {User request}
###Data files: {All data files}
Please solve the request by Python Code.
Step-by-Step ReAct
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and generate your thoughts and actions for each step, and then finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please summarize the results of the code execution just now and think about whether this result accomplishes what you planned for this step.
### {Action: Yes, I observed that this function successfully fetched the data...}
Step-by-Step Reflexion
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and then generate your thoughts and actions for each step. After the execution of your current action, you need to reflect on the results until your current plan has been successfully completed. Then you think about the next step and then generate your next action, and finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please record the compilers return results just now and think about whether this result accomplishes what you planned for this step.
### {Action: No, I observe that the compiler returns an error...}
### Reflection Prompt: Please reflect on the error returned by the compiler and regenerate a new Python code to resolve the issue. If the compilation passes without any errors, reflect on whether the current result is what you planned to do.
### {Action: I revise my solution as follows: def get-data2()....}
Multi-agent collaboration
###Instruction For Manager: You are the manager of the project team and you need to lead your team to fulfill user requests. You have two experienced programmers under you (Programmer-A and -B) and you need to assign them the same or different tasks according to the user request, then organize the discussion, and finally solve the problem.
###Instruction For Agent1/Agent2: You are an experienced programmer. Your team has a colleague who is also a programmer and a manager. You need to write code according to the managers arrangement, discuss it with them, improve your program, and get a consensus conclusion.
###User Request: {User request}
###Data files: {All data files}
-------- Phase1 Discussion for Task Assignment -------
###Task Assignment Prompt For 3 Agents: Now lets start discussing how to solve user problems. The speaking order is manager, A and B. After a round of discussion, the manager decides on a preliminary solution and assigns specific tasks to the two programmers
###Manager Response: {To fulfill the request...}
###Programmer A Response: {Yes, I agree with the manager. We should first,....}
###Programmer B Response: {That’s a good solution. We also should...}
###Manager Response: {After discussion, Programmer A should..., B should ....}
-------- Phase2 Coding and Discussion-------
###Prompts For 3 Agents: Two programmers complete their respective tasks as assigned by their manager. Then the manager organizes a discussion on whether the two options and their running results are reasonable. Lastly, two programmers need to regenerate the code.
###Programmer A Response: {Solution1 Generated by A}
###Programmer B Response: {Solution1 Generated by B}
###Prompts For 3 Agents: Please evaluate two solutions. {Solution1, Result1, Solution2, Result2}
###Programmer A Evaluation: {I think ...}
###Programmer B Evaluation: {I agree with ...}
###Programmer A Regeneration: {Revised Solution1}
###Programmer B Regeneration: {Revised Solution2}
-------- Phase3 Discussion-------
-------- Phase4 Final Decision-------
###Prompts For manager: Please output a complete and correct solving code for user request based on the previous discussions of the group.
###Manager Decision: {Final Solution}
Appendix G Case Study
Detailed cases of data exploration and workflow deployment are presented in Sections G.1 and G.2.
G.1 Cases for Data Exploration
---- The First Request: Interface Design -----
Input Request: Get the daily opening price data of Bank of Hangzhou from Jan 23, 2018 to Mar 13, 2019 and plot the trend.
Response: First, we need to design a function to get the opening price data from the banks database. Then, we need another function to process this data and prepare them for graphical representation. Finally, we need a function to plot the trend chart.
New Design Interfaces:
{Interface1=Name:getBankData(), Function:.., Input:.., Output:..}
{Interface2=Name:prepareDataForPlotting(), Function:.., Input:.., Output:..}
{Interface3=Name:plotData(), Function:.., Input:.., Output:..}
The solving process for request: These three new interface functions will meet your needs. In your program, you would first get the data using getBankData, then process the data using prepareDataForPlotting, and finally use plotData to plot the trend.
Complete code: {Complete code using three new interfaces...}
---- The Second Request: Interface Definition Phase-------
Input Request: Get the monthly closing price data of Ningde Times from January 23, 2018 to March 13, 2019, and plot the trend.
Response: To fulfill this requirement, we design a new interface function. In addition, I need to use two existing interfaces to fulfill your requirement.
New Designed Interfaces:
{Interface4=Name: getStockData()..}
Existing Interfaces
{Interface1=Name:getBankData(), ...}
{Interface2=Name:prepareDataForPlotting(), ...}
{Interface3=Name:plotData(), ...}
The solving process for request: You need to call these functions to meet your needs. First, use new interface {getStockData()} to get the data of Ningde Times, then use existing interface {prepareDataForPlotting()} to process this data, and finally use {plotData()} to plot the trend...
Complete code: {Complete code using three new interfaces...}
---- Interface Optimization Phase -------
The reason for merging: Yes, these interface functions can be merged and reformulated to eliminate duplicate functions. We can merge getBankData() and getStockData() into a more generic function get-financial-data(), which can fetch any financial data, including bank and stock data. In addition, we can keep the prepareDataForPlotting and plotData functions, as they can also be applied to different types of financial data]
{Interface1=Interface Name:get-financial-data(), ...}
{Interface2=Interface Name:prepareDataForPlotting(), ...}
{Interface3=Interface Name:plotData(), ...}
G.2 Cases for Real-time Request
### Request: Today is 2019-03-13. The user request is to please help me plot the stock price trend of Guizhou Maotai from January 23, 2018 to today.
### Rewritten Request: Please help me plot the closing price of Guizhou Maotai from Jan 23, 2018 to Mar 13, 2019 (Time:20180123-20190313, Location: China, Object: Guizhou Maotai stock price, Format: Line).
###Plan: task1={"stock_task": "Get the time series data of Guizhou Maotais stock price from 20180123 to 20190313"}, task2={"visualization_task": "Plot the stock price of Guizhou Maotai from 20180123 to 20190313 on a line chart"}
###Function Call:
step1={
"arg1": ["Guizhou Maotai", "20180123", "20190313", "daily"],
"function1": "get_stock_prices_data",
"output1": "result1",
"description1": "Guizhou Maotai daily stock price (open, close, high, low, etc.) time series data"},
step2={
"arg1": ["result1","close"],
"function1": "calculate_stock_index",
"output1": "result2v,
"description1": "Guizhou Maotais closing price time series data"},
step3={
"arg": ["result2", null, "line", "Guizhou Maotai closing price chart from January 23, 2018 to March 13, 2019"], "function": "plot_stock_data",
"output": "result1",
"description": "Guizhou Maotai Stock Price Line Chart"
Figure G2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
Figure G3: We have listed some of Data-Copilot’s self-design interfaces, including five categories. Most of the interfaces are a combination of several simple interfaces. For example, the text marked in blue indicates that the interface includes multiple functions.
Figure G4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
Figure G5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
Figure G6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
Figure G7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
Figure G8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The contribution of the article is clearly stated in Abstract and Section 1 (Introduction), and is fully consistent with the experiment (Section 3).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Appendix Section C.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not deal with the theoretical derivation part.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have provided all the experimental details in the article (Section 4.1), prompt design (Appendix A, B, F.1 F.2 F.3).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have provided a code implementation for our system. The codebase has been open-sourced.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided all the experimental details in the article (Section 4.1), prompt design (Appendix A, B, F.1 F.2 F.3).
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide details of the experiment, reporting the mean and standard deviation after multiple experiments in Table 2.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In our experiments, we report the model of the API interfaces used and the tokens consumed for each experiment.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [N/A]
  44. Justification: This paper does not involve any issues related to the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We propose an autonomous data analysis agent that does not generate any negative social impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: Our method does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We utilized financial data provided by Tushare, which has been duly cited in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: The paper does not release new assets
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We invited experts to annotate and evaluate the data, providing them with detailed instructions and guidelines.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not present these risks.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v7

@@
 Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
-To achieve this, as shown in Figure 1, Data-Copilot contains three steps for data exploration:
+As shown in Figure 1, Data-Copilot contains three key steps for data exploration:
   1. 1.
-    Request Exploration: Data-Copilot explores potential requests and mines a broader range of human need scenarios from data.
-It involves a "self-instruct" process to generate massive simulated requests based on all data schemas (types and formats) and seed requests.
+    Explore data and synthesize requests: Data-Copilot discovers potential requests and a broader range of human need scenarios from data.
+It involves a "self-exploration" process to generate massive synthesized requests based on all data schemas and seed requests.
   2. 2.
-    Interface Design: Our agent generates code for each request and then considers whether these requests have similar demands and whether general interfaces can be abstracted.
+    Interface design and test: Data-Copilot generates modular code for request solving. Each interface is a functional module, and test cases are generated to verify the designed interface.
   3. 3.
-    Interface Optimization: To improve the quality of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces by reflecting on feedback from the compiler.
+    Interface optimization: To improve the versatility of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces based on the compiler’s feedback.
 After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks.
 When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests.
-For different request scenarios, Data-Copilot can flexibly design different interface invocation procedures, such as serial workflows, i.e., step-by-step invocation, parallel workflows, and loop workflows.
+For different requests, Data-Copilot can flexibly deploy various invocation structures, such as step-by-step serial workflows, parallel, or loop.
+It can even output a hybrid form of interface workflows and raw code for some "unfamiliar" requests.
 Our contributions are threefold:
 Figure 2: Overview of Data-Copilot.
-Data Exploration: It devises a self-request process to generate numerous potential requests based on data schemas.
-Then it interleaves to design and optimize the interfaces to solve these requests.
-The whole exploration-design-optimization process is operated in advance, generating many general and error-free interface modules, such as data retrieval, computation, and visualization.
-Workflow Deployment: Upon receiving real-time requests, LLM invokes these interfaces, forming an interface workflow with different structures for request solving.
+Data Exploration: First, it performs a self-exploration process to uncover potential human requests from data sources.
+Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces one by one.
+This exploration-design-optimization process is operated in advance.
+Workflow Deployment: Upon receiving real-time requests, Data-Copilot can quickly invoke existing interfaces and deploy a workflow for familiar requests, or flexibly combine interfaces and raw code for "uncovered" requests.
   1. 1.
@@
 Data-Copilot operates in two phases: Data Exploration and Workflow Deployment.
-In the first phase (Section 2.1), Data-Copilot designs a self-request process to discover numerous potential needs based on data schemas (Step 1: exploration).
-Then based on these "self-instruct" requests, it generates code and abstracts universal interface modules for each request (Step 2: design).
-After that, it optimizes similar interfaces and reflects on the code compiler feedback, ensuring each interface is universal and correct (Step 3: optimization).
-Finally, Data-Copilot acquires many generic, error-free interfaces for subsequent use.
-The whole exploration-design-optimization process is operated in advance.
-When faced with real-time requests (Section 2.2), Data-Copilot adeptly comprehends user intentions, and invokes pre-designed interfaces for request solving, significantly reducing the need for tedious labor.
+In the first phase (Section 2.1), Data-Copilot designs a self-exploration process to discover numerous potential needs based on data schemas (Step 1: exploration).
+Then based on these synthesized requests, it abstracts universal interface modules for each request (Step 2: design).
+After that, it optimizes similar interfaces and tests their correctness, ensuring each interface is correct and universal (Step 3: testing and optimization).
+The whole exploration-design-optimization process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
+When faced with real-time requests (Section 2.2), Data-Copilot can flexibly invoke pre-designed interfaces or directly generate raw code based for the user’s request.
+In most cases, existing interfaces can cover the majority of real-world requests, significantly enhancing both success rate and response speed.
 We provide a detailed prompt for two phases in Section F.1 and F.2.
@@
 Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability.
-Similar to humans, Data-Copilot aims to explore data and derive insights from vast data sources.
-These insights include understanding the relations between the data, discovering the potential requests associated with the data, and organizing the code framework efficiently.
-Based on this, Data-Copilot summarizes these exploratory insights into many reusable code components and optimizes its performance.
-It significantly reduces the code complexity for real-time responses, thus improving the success rate.
-Request Exploration
-To explore data and mine insights, we design a self-request process.
+Similar to this, Data-Copilot also autonomously explores data and derives insights from vast data sources, including the relations between the data, and the potential requests associated with the data.
+Then Data-Copilot abstracts these exploratory insights into numerous reusable code components (interfaces), testing their correctness and optimizing their generality.
+This process of exploring data, identifying common requests, designing general interfaces, and testing and optimizing their performance is conducted in advance on its own.
+Self-Exploration Request
+To explore data and mine insights, we design a self-exploration process.
 Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario.
-This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs must generate requests specifically related to these data.
+This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs should generate requests specifically based on provided data.
 As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years.
@@
 First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
-###Parsing file for Stock_data:{...},
+###Parsing file for Stock_data:{...}, ....
 How to ensure the quality of the generated requests?
@@
 Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests.
-Upon manually check some requests, we find that the generated requests generally met our expectations.
+Upon manually checking some requests, we found that the generated requests generally met our expectations.
 Interface Design
-After generating massive requests through the self-request, Data-Copilot considers abstracting universal interfaces from these requests.
-First, we need to define what an interface is in our paper.
-Similar to human programming practices, we refer to a function module abstracted from the code as an interface.
-Each interface is composed of a name, parameters, a functional description, and the implementation code, performing specific functions such as data retrieval, computing, and plotting.
-To abstract general interfaces from all requests, we design an interwoven process, performing interface design and interface optimization alternately for each request.
-Firstly, we feed related data parsing files and existing interfaces stored in the interface library (empty at first iteration) into LLMs as a prompt.
-LLMs need to generate a complete code for each request.
-When generating, LLMs are instructed to prioritize using the existing interfaces from the previous iterations.
-If these interfaces are insufficient to address the request, the LLMs design a new interface.
+After generating massive requests through self-exploration, Data-Copilot abstracts many universal interfaces from these requests.
+First, we have to clarify what an interface is in our paper.
+Similar to human-defined functions, an interface is a code module consisting of a name, parameters, a functional description, and an implementation code.
+It performs specific tasks such as data retrieval, computation, and visualization.
+In our paper, each interface is designed, tested, and optimized iteratively by Data-Copilot.
+First, starting from the initial request, we iteratively feed self-exploration requests and related data parsing files into LLMs, prompting the LLM to design complete code modules for request solving.
+Each code module corresponds to an interface and is stored in the interface library.
+During each iteration, LLMs are instructed to prioritize utilizing the existing interfaces within the library.
+If the available interfaces are insufficient for request solving, the LLMs design a new interface.
 As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...).
 As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
+Interface Testing
+After designing a new interface for a request (request-i $\rightarrow$ interface-i), Data-Copilot autonomously tests its correctness based on compiler feedback.
+First, Data-Copilot uses the request-i as the seed to generate massive similar requests as test cases.
+Then it tests the new interface-i one by one.
+E.g., in Fig.
+2, for newly designed interface-i: Query-GDP(), Data-Copilot mimics request-1 to generate many similar requests to test the Query-GDP(): "I want to see USA’s GDP over the past 10 years", "I want to see China’s GDP for the last year".
+If the interface passes all test cases, it is retained; otherwise, Data-Copilot utilizes compiler error feedback for self-reflection, correcting the erroneous code snippets until it successfully passes the tests.
 Interface Optimization
-To improve the quality of interfaces, Data-Copilot optimizes them in two ways: merging similar interfaces to make them more universal and reflecting on compiler feedback to reduce errors.
-This optimization process is intertwined with the interface design for each request.
-Firstly, similar to software developers wrapping similar code modules to reduce redundancy, Data-Copilot considers whether each newly designed interface can be merged with the previous interfaces.
-LLMs are instructed to check the functionality, parameters, and logic of the new and existing interfaces, and merge those similar interface modules to create a new one.
+To further improve the generality of interfaces, Data-Copilot also optimizes similar interfaces.
+Similar to how human software developers, each time a new interface is designed, this optimization process is triggered: evaluating whether the newly designed interface can be merged with previous ones.
+  • Retrieve similar interfaces.
+After a new interface is designed, we first retrieve the top N similar interfaces from the existing interface library.
+Specifically, we use gte-Qwen1.5-7B-instruct (Li et al., 2023d) to obtain embeddings of each interface code and then calculate their similarity.
+  • Decide whether to merge.
+Then LLMs are prompted to evaluate the functionality, parameters, and logic of the new interface and the top N similar interfaces, autonomously deciding whether to merge the new interface with existing ones.
+If the LLM deems no merging necessary, the new interface is retained.
 As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...).
-This interface optimization process makes each interface more general and rational.
-Besides, to reduce code hallucinations or formatting issues, Data-Copilot also self-reflects on the feedback of code execution for each request if compilation errors occur.
-We have LLMs perform self-reflection until the code passes or the maximum retry number is reached.
-This process ensures the correctness and usability of each interface.
-For each request, Data-Copilot alternates the above two steps: interface design and optimization until all the requests can be satisfied by these interfaces.
-Finally, as shown in Figure G3, Data-Copilot produces five types of interfaces: data acquisition, processing, prediction, visualization, and DataFrame manipulation.
+This optimization process makes each interface more general and unique, reducing interface redundancy.
+  • Test the optimized interface.
+When two interfaces are merged, Data-Copilot also needs to test the newly merged interface.
+The test cases from the two original interfaces are used to validate the merged interface.
+The output of the new interface must be consistent with the two original interfaces in both format and content.
+For each request, Data-Copilot alternates the above three steps: interface design, testing, and optimization until all the requests can be covered by these interfaces.
+As shown in Figure G3, Data-Copilot designs many interfaces for different task and operation types, e.g., data acquisition, prediction, visualization, and DataFrame manipulation.
 We provide some detailed cases in Section G.1.
-Besides, to handle newly added data sources, Data-Copilot can repeat the aforementioned exploration-design-optimization process on new data.
-It can create some new interface modules for these new data.
 2.2 Workflow Deployment
-In the first phase, Data-Copilot designs many generic interfaces with different functionalities.
-They are versatile and easy to use, simplifying the code generation of real-time responses.
-As shown in Figure 2, when faced with human requests, Data-Copilot only needs to invoke related interfaces and form an interface workflow.
-A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex call flow diagram.
-This workflow transforms raw data into visualized results.
-The detailed prompts and cases are shown in Sections F.2 and G.2.
+In the first phase, Data-Copilot designs many generic interfaces with different functionalities in advance.
+They are versatile and error-free, capable of handling the majority of human requests in real-world scenarios.
+As shown in Figure 2, Data-Copilot only needs to invoke the relevant interfaces for human requests, which is more accurate and efficient (interface workflow).
+Furthermore, for some long-tail or new requests that cannot be covered by the existing interfaces, Data-Copilot can flexibly generate raw code to handle them (interface-code hybrid strategy) and document these requests for updating the interface library in the future.
+Task Selection and Interfaces Retrieval
+Considering the significant differences between the interfaces involved in different requests, it is unnecessary to load all interfaces every time.
+We design a simple yet efficient interface retrieval strategy: hierarchical retrieval.
+Specifically, we organize all designed interfaces in a hierarchical structure.
+Each interface is grouped into different tasks (stock task, fund task, etc.) and different operation types (data acquisition, processing, visualization, etc.).
+Upon receiving a real-time request, Data-Copilot first determines the appropriate task types and required operation types and then loads the interfaces associated with these types for subsequent workflow planning.
+This can significantly reduce the number of interfaces in the prompt.
+Planning Workflow with Different Structures
+After reading the descriptions of all retrieved interfaces, Data-Copilot plans an interface workflow to process the user request.
+A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex loop structures.
+Therefore, Data-Copilot needs to autonomously determine which interfaces to invoke, in what sequence, and which parameters to populate.
+The interface invocation workflow is output in JSON format: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}.
+The prompts and cases are shown in Sections F.2 and G.2.
+As shown in Figure 2, Data-Copilot designs interface workflows with different structures, e.g., sequential, parallel, or loop.
+For instance, the first request in Figure 2: "Compare the return of CSI-300, GEM and CSI-1000 this year", Data-Copilot plans a parallel workflow with 5 steps: Data Acquisition(), Calcuate Return() in parallel for calculating three indices, followed by Visualization().
+In the second case, it deploys a loop workflow by calling the Loop().
+Interface-Code Hybrid Generation
+In the real world, it is inevitable to encounter "uncovered" requests that cannot be addressed by existing interfaces.
+As a remedy, Data-Copilot adopts an interface-code hybrid generation strategy.
+Specifically, Data-Copilot prioritizes invoking existing interfaces to resolve user requests.
+It can solve most common user requests.
+However, if the deployed workflow continues to fail, or if Data-Copilot proactively determines that the current request cannot be resolved by existing interfaces, our system is allowed to generate raw code directly or generate a combination of raw code and interfaces for request solving.
+This design endows Data-Copilot with the flexibility to handle diverse requests.
+Besides, each time Data-Copilot encounters such "uncovered" requests, it also records them in the document.
+After accumulating a certain number of new requests, we re-initiated the Data Exploration stage (Section 2.1) to develop new interfaces for these newly emerged requests.
+It empowers Data-Copilot to continuously update its interface library throughout human interactions.
+Table 1: Statistics on human-proposed requests and self-exploration dataset.
+Source #Cases Form Usage Ratio of 4 Tasks Ratio of 3 Complexities
+Human-proposed 173 Query Seed Set 8: 4: 4: 1 2.0: 1.5: 1
+Self-Exploration 3000 Query Interface Design 8.7: 3.3: 4.2: 1 1.9: 1.5: 1
+Self-Exploration & Human 547 Query, Label Test Set 8.6: 3.3: 4.3: 1 1.8: 1.4: 1
+3 Qualitative Evaluation
+3.1 Environment and Data Sources
+Data-Copilot is evaluating on Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data.
+Similar to many real-world financial companies, Data-Copilot utilizes data interfaces provided by Tushare [^1] https://tushare.pro/ to access vast amounts of financial data, including time-series data spanning over 20 years for more than 4,000 stocks, funds, and futures.
+In the first phase, we use a strong LLM (e.g., gpt-4-turbo) for data exploration and a lightweight LLM (e.g., gpt-3.5-turbo) for workflow deployment.
+3.2 The Creation of Dataset
+Human-proposed Request as Seed.
+We invite 10 students in economics to submit 50 requests each.
+These requests cover a wide range of common needs, including stock tasks, fund tasks, and corporate tasks, and also involve different difficulty levels.
+We treat these 500 human-proposed requests as a distribution of real-world scenarios, where the ratio of four task types is Stock: Fund: Corporation: Other = 8: 4: 4: 1.
+Then we filter out highly similar requests by computing their sentence embedding.
+Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
+Self-Exploration Request Set.
+As mentioned in Section 2.1, we employ self-exploration to expand the request set.
+Specifically, we feed 173 human-proposed seed requests along with all data descriptions and schema into GPT-4 for data exploration and request synthesis.
+When self-exploration, we use keywords to control the distribution of the synthesized requests, ensuring that they can closely align with real-world scenarios.
+In this phase, we generate 6480 requests.
+After that, we filter out highly similar requests and retain 3547 requests.
+Then we adopted a stratified sampling strategy to sample 547 instances as the test set from each task type.
+The remaining 3000 requests are used for interface design.
+We analyze the number of subsets categorized by four task types (stock, fund, corporation, and others) and three complexity levels (single-entity, multi-entities, and multi-entities with complex relations).
+The results indicate that our synthesized requests closely align with human distribution.
+The detailed statistics of three datasets are shown in Tables 1, B1 and B1.
+Table 2: Human evaluation on the human-proposed seed set and self-exploration requests.
+Number Task Difficulty Request Rationality Expression Ambiguity Answer Accuracy
+Seed Set 173 3.5 4.3 4.4 -
+Test Set 547 3.9 4.2 4.0 4.8
+Annotate Answer Table for Test Set.
+Human annotators are instructed to annotate data-table as answers for 547 test requests, which is used for benchmarking our system’s output.
+Specifically, annotators manually retrieve and process the corresponding data based on the testing request, and record the final data table before chart plotting.
+These data tables are used as answer annotations.
+For example, for the request: "I want to see China’s GDP trend over the past 5 years", the answer data-table is "[2023: 17.8 trillion, 2022: 17.9 trillion, 2021: 17.8 trillion, 2020: 14.6 trillion, 2019: 14.3 trillion]".
+The Quality of Self-exploration Requests.
+To assess the quality of self-exploration requests, we invited four additional graduate students to manually evaluate our test set (546 requests) and human-proposed requests (173 seed set) on four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy.
+We provide detailed guidance for each criterion in Section B.2.
+The results of two sets are shown in Table 2.
+We observed that the synthesized requests exhibit a comparable quality to human-proposed requests, with slightly higher difficulty, ambiguity, and similar rationality.
+It ensures that our test set can reflect most of the real-world demands.
+Besides, the evaluators gave a high score of 4.8 on the label (answer table) of our test set, which also ensures the accuracy of our dataset.
+3.3 Qualitative Results
+The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request.
+The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces.
+For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
+Data-Copilot designs versatile interfaces through data exploration.
+After exploring the Chinese financial data and processing 3000 self-exploration requests, Data-Copilot designed a total of 73 interfaces, comprising 5 functionalities and 4 tasks: data acquisition, index calculation, table manipulation, visualization, and general processing logic.
+These tasks include stocks, funds, corporations, and others.
+We present some representative interfaces and the statistics of the interfaces in Figure G3.
+Most interfaces are highly versatile with multiple functions.
+For example, get-stock-prices-data() interface can retrieve data on a daily, weekly, or monthly basis.
+These versatile interfaces endow the Data-Copilot with the capability to address complex requests.
+Data-Copilot can flexibly deploy interfaces or raw code for diverse demands.
+As mentioned in  Section 2.2, for some "new" requests, Data-Copilot is capable of addressing them by combining both raw code and existing interfaces.
+As shown in Figure 2 (3rd example), the user requests a radar chart, which was not directly supported by the available interface.
+Data-Copilot first leverages the existing interface to retrieve and process the data.
+Then based on the dataframe produced by the last interface, it generates raw code to render a radar chart.
+This design grants Data-Copilot greater flexibility to handle the constantly emerging new requests.
+Interface Workflow is human-friendly, offering higher interpretability.
+Upon executing the workflow via the compiler, Data-Copilot produces the visualized results.
+Also, it provides a comprehensive summary of the entire processing process based on the workflow.
+As shown in Figures 2, G2, G4, G6, G5 and G7, Data-Copilot’s outputs (workflow, visualized results and summary) are highly intuitive and clear.
+It encapsulates complex code processing logic into a step-by-step interface invocation workflow, which is easier for humans to understand and inspect.
+Table 3: Pass@1 of Data-Copilot and direct code generation (Raw Code) when using different LLMs for workflow deployment: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Llama-3-70B-chat, Llama-2-70B-chat, Llama-2-13B-chat, Codellama-13B, Vicuna-13B-v1.5).
+GPT-4o GPT-4 GPT-3.5 Llama3-70B Llama2-70B Llama2-13B Codellama13B Vicuna
+Raw Code 73.6 70.0 28.5 52.2 29.6 17.1 21.0 17.3
+Ours 77.9 $\uparrow$4.3 74.6 $\uparrow$4.6 70.2 $\uparrow$41.7 70.9 $\uparrow$18.7 49.3 $\uparrow$19.7 36.9 $\uparrow$19.8 50.8 $\uparrow$29.2 33.2$\uparrow$15
+4 Quantitative Evaluation
+4.1 Experiments Settings
+Baselines and Experiments Details We first compare Data-Copilot with direct code generation (denoted as Direct-Code) using different LLMs.
+Besides, we also compare Data-Copilot with many prompting and agent strategies, including: 1.
+ReAct (Yao et al., 2022): LLMs use neural language (thought) and code to resolve the user request step-by-step.
+2.
+Reflexion (Shinn et al., 2023): LLM can reflect on the feedback from the compiler at each step, progressively fulfilling user requests.
+For the Reflexion strategy, we limit it to a maximum of two rounds of reflection.
+3.
+Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion.
+For a fair comparison, Data-Copilot employs the same LLM as all baselines for evaluation when invoking pre-designed interfaces.
+The detailed prompts of all baselines are shown in Section F.3.
+Evaluation Methods.
+All methods are evaluated on the 547 testing instances.
+We focus on the accuracy of the final results and the efficiency of the response.
+Considering that the output of most requests is images or charts, we evaluate from two perspectives: whether the data used for plotting is correct and whether the drawn image meets human requirements.
+We use GPT-4-turbo and GPT-4o to conduct automatic evaluations.
+(1) Data-table Evaluation: We first obtain the dataframe before visualization, then use GPT-4-turbo to compare this data table with the answer table (labeled by humans in Section 3.2).
+(2) Image Evaluation: To evaluate whether the final image meets human requests, we feed the human request, human-annotated answer table, and the output image into GPT-4o for evaluation.
+We instruct GPT-4o to check whether visual elements in the image meet the human requests, including numerical points, lines, axes, image aesthetics, and style.
+Finally, the results of data and image evaluation are combined as Pass@1.
+Lastly, we calculate the total number of tokens consumed by each method (denoted as #Token), which represents the response efficiency.
+Table 4: We demonstrate the accuracy and efficiency of all methods with gpt-3.5-turbo
+Methods Pass@1(%) #Token
+Direct-Code 28.5 $\pm$2.1 823
+ReAct 44.1 $\pm$3.4 1515
+Reflexion 59.0 $\pm$3.9 2463
+Multi-Agent 57.4 $\pm$1.8 2835
+Data-Copilot 70.2 $\pm$2.5 561.2
+Data-Copilot + ReAct 71.5 $\pm$1.7 834
+Data-Copilot + Reflexion 71.8 $\pm$2.2 978
+Table 5: All samples are divided into 3 complexity levels. We compute Pass@1.
+Methods Single Multiple Complex Overall
+Direct-Code 42.4% 29.0% 1.6% 28.5%
+ReAct 56.3% 48.6% 14.6% 44.1%
+Reflexion 67.8% 55.1% 48.1% 59.0%
+Multi-agent 63.2% 64.7% 35.5% 57.4%
+Data-Copilot 71.8% 70.0% 67.1% 70.2%
+4.2 Comparison Results
+Data-Copilot Significantly Reduces the Risk of Failure When Real-time Deployment.
+As demonstrated in Table 3, Compared to directly generating code, Data-Copilot achieves significant improvements across all LLMs, with particularly strong gains of +41.7, +18.7, and +29.2 on GPT-3.5-turbo, Llama3-70B, and Codellama-13B, respectively.
+By decoupling direct code generation into two stages: interface design in advance and real-time workflow deployment, Data-Copilot can elevate many smaller LLMs (e.g., Lllama3-70b, GPT-3.5) to GPT-4 level during real-time deployment.
+Through careful comparison, we observe that these pre-designed interfaces significantly reduce the occurrence of errors during real-time responses, as they have already been optimized and validated during the design phase.
+In contrast, directly generating code from scratch makes it easy to overlook certain details or make mistakes.
+Besides, we also found that even with a powerful LLM like GPT-4, Data-Copilot also achieves an improvement of over +4%, validating its generalizability.
+Data-Copilot Outperforms Advanced Agent Strategies in Both Accuracy and Efficiency.
+As shown in Table 5, Data-Copilot surpasses all baseline strategies in success rate (Pass@1), and efficiency (#token).
+Compared to the best baseline (Reflexion), Data-Copilot achieved an +11.2% improvement in accuracy and a -75% reduction in token consumption.
+Besides, as shown in the last two rows of Table 5, Data-Copilot can seamlessly integrate with agent strategies such as ReAct and Reflexion into the workflow deployment.
+For example, when combined with ReAct, Data-Copilot invokes an interface, obtains an intermediate result, and then reasons to invoke the next interface, further improving its performance by +1.3.
+Data-Copilot Reduces Repetitive Generation in the Real World.
+In the real world, numerous requests are similar or even repetitive.
+Data-Copilot can save 70% of the baseline’s consumption because its output is very concise and only contains interface names and arguments.
+In contrast, baseline strategies have to repetitively generate complete code for each request, consuming many more tokens.
+Table 6: Different LLM combinations for two stages. Using a weaker LLM for interface design affects both the interface’s quantity and quality.
+Direct Code Generation: GPT-3.5: 28.3, GPT-4: 70
+LLM for Workflow Deployment
+GPT-3.5 GPT-4
+LLM for Interface Design GPT-3.5 31.9 49.6
+GPT-4 70.2 74.6
+Table 7: We ablated 3 designs from Data-Copilot.
+Methods Pass@1(%)
+Direct Code Generation 28.5
+Data-Copilot 70.2
+$w/o$ Self-Exploration Request 35.9
+$w/o$ Interface Optimization 42.1
+$w/o$ Interface-Code Hybrid 63.8
+4.3 Analysis: Why Data-Copilot Can Bring Improvements?
+Data-Copilot Exhibits a Pronounced Advantage on More Complex Scenarios.
+To better analyze the behavior of Data-Copilot, we categorize each test request into different subsets based on the number of entities it involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations).
+The statistics for each subset are shown in Section B.1.
+As shown in Table 5, for the single entity subset (denoted as Single), which represents the simplest case, the baselines and our method perform similarly, with only a 4% improvement.
+However, for the more challenging multi-entities (denoted as Multiple) or multi-entities with complex relation (denoted as Complex), our improvements are substantial, with enhancements of 5.3% and 19%, respectively.
+We find that in these complex cases, baseline methods often generate logically incorrect code on the first attempt and may omit some critical steps.
+In contrast, Data-Copilot only needs to invoke versatile interfaces and deploy workflows, significantly reducing the complexity of real-time responses.
+These findings highlight the advantages of Data-Copilot, which reduces task complexity by pre-designing interfaces in advance.
+Interface Quality is the Key Factor.
+We analyze the effects of using different LLM combinations for two stages: interface design and workflow deployment.
+As shown in Table 7, we observed that when using a weaker LLM for interface design, the effectiveness is significantly diminished.
+For example, using GPT-3.5-turbo (1st stage) and GPT-4-turbo (2nd stage) resulted in a score of only 49.6, which is even lower than directly using GPT-4-turbo for code generation (70.2).
+This phenomenon can also be seen in other combinations.
+After manually checking the interfaces in this case, we found these interfaces designed by GPT-3.5-turbo are limited in number and of poor quality.
+Therefore, we chose to use GPT-4-turbo for interface design and optimization, while employing GPT-3.5-turbo for real-time deployment, achieving a balance between accuracy and efficiency.
+4.4 Ablation Study
+As shown in Table 7, we ablate Data-Copilot from the following three aspects:
+  1. 1.
+    Self-exploration Request.
+We ablate the self-exploration request process Section 2.1.
+Instead, we directly used seed requests for interface design and optimization.
+In this case, the performance dropped significantly by 34.3.
+This may be because there are too few seed requests, which are insufficient to design universe interfaces.
+  2. 2.
+    Interface Optimization.
+We ablate the interface optimization phase during interface design.
+It means every successfully designed interface was retained in the interface library.
+We observed that performance was also significantly affected (-28 points).
+Without interface optimization, there are many similar interfaces, which could hurt the effect of workflow invocation.
+  3. 3.
+    Interface-Code Hybrid Generation.
+Data-Copilot only can invoke interfaces during workflow deployment.
+Directly generating code is not allowed even when encountering requests that cannot be resolved by interfaces.
+After ablation, we observe a performance drop of 6.4 points.
+These losses are mainly caused by new requests that cannot be covered by existing interfaces.
+It highlights the flexibility of Data-Copilot, which can flexibly solve familiar requests by invoking interfaces or handle new requests by generating code directly.
+5 Conclusion
+We propose a data analysis agent, Data-Copilot, to address a wide array of data-related tasks.
+Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests.
+It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
+Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
+Appendix
+Appendix A Preprocessing of Workflow Deployment
 Intent Analysis
@@
 To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
-Planning Workflow
-After providing the user request and a description of the relevant interfaces, Data-Copilot plans a logical interface workflow to process the user request.
-Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot.
-The LLMs are instructed to output in JSON format for each step invocation: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}.
-For different requests, Data-Copilot designs interface workflows with different structures, such as sequential, parallel, or loop.
-For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot plans a five-step workflow.
-In the first two steps, it dispatches Data Acquisition(), Calcuate Return() in parallel for three indices.
-The last three steps successively invoke the Visualization() to sequentially plot three indices.
-In the second case, it deploys a loop workflow by calling the Loop().
 Multi-form Output
@@
 3) Visualizing the output.
-Table 1: Left: Statistics on human-proposed requests and self-request dataset. Right: Detail Statistics on different request types and workflow structures.
-Source #Cases Form Usage
-Human-proposed 173 Query Seed
-Self-Request 3000 Query Interface Design
-Self-Request + Human-annotated 547 Query, Ans Testing
-Request Types Workflow Types
-Sequential Parallel Complex Overall
+Appendix B Evaluation Details
+B.1 The Definitions of Three Complexity Levels for Testset
+We categorize our test set into three types based on the number of entities involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations).
+The ststics of three subset are shown in Table B1.
+  • Single entity: Requests involve a single entity and can be resolved step-by-step. E.g., "query based on a specific condition".
+  • Multiple entities: Requests require processing multiple entities simultaneously. E.g., "compare a certain metric across multiple entities".
+  • Multiple entities with complex relations: Involving multiple entities and containing loops, nesting, and other intricate logic.
+E.g., "List the top 10 stocks by yesterday’s price increase that also in the internet industry."
+Table B1: Statistics of four task types and three complexity levels on our test set.
+Task Types Request Complexity
+Single Entity Multiple Entities Multi-entities with Complex Relation Overall
 Stock 79 106 87 272
@@
 Total 237 184 126 547
-3 Qualitative Evaluation
-3.1 Environment and Data Sources
-We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data.
-To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies.
-The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance.
-In the first phase, we use the gpt-4-0613 for data exploration design and gpt-3.5-turbo-0613 for workflow deployment.
-3.2 Seeds and Datasets
-Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each.
-These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate.
-2) involves different levels of difficulty.
-Then we filter out highly similar requests by computing the sentence embedding of each request.
-Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
-en Self-request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request to expand the request set.
-Specifically, we feed 173 human-proposed requests along with all data descriptions and schema into GPT-4.
-We instruct GPT-4 to generate massive requests.
-In this phase, we generate 6480 requests, then compute the semantic similarity between all requests, and filter out those with high similarity, and retain 3547 requests.
-We randomly select 3000 requests from the total set for subsequent interface design.
-Meanwhile, the remaining 547 requests are treated as a test set.
-We manually annotate the final answers, task types, and workflow types for each testing instance.
-The statistics of three datasets are shown in Tables 1 and B1.
-3.3 Qualitative Results
-The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request.
-The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces.
-For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
-Data-Copilot designs versatile interfaces through data exploration.
-In Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, manipulation, and visualization.
-These diverse interfaces endow the LLM with the capability to address complex requests.
-For instance, get-stock-prices-data() interface integrates the functions of acquiring data by day, week, and month, controlled by argument: freq, which is similar to the process of programmer development.
-Data-Copilot deploys workflows with different structures.
-Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests.
-As shown in Figure 2, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface.
-It iteratively calculates the financial indicators for each stock.
-Please refer to Appendix E for more visualization cases.
-4 Quantitative Evaluation
-We quantitatively assess Data-Copilot in processing real-time user requests, where it could invoke the interfaces designed in the first phase.
-4.1 Experiments Settings
-Baselines 1.
-Vanilla LLM (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests.
-2.
-ReAct (Yao et al., 2022): The LLM resolves the user request using neural language and code step-by-step.
-3.
-Reflexion (Shinn et al., 2023): LLM can also reflect on the feedback from the compiler at each step, progressively fulfilling user requests.
-For the Reflexion strategy, we limit it to a maximum of two rounds of reflection.
-4.
-Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion.
-All methods are evaluated on the 547 testing instances (Section 3.2) three times.
-The detailed prompts of all baselines are shown in Section F.3.
-Metrics We design three metrics for solution code and answer: Exec@1 Pass@1 and Efficiency.
-Exec@1 evaluates whether the code can be executed by a compiler without errors.
-Pass@1 measures the correctness of the final answer.
-Considering the diverse output formats, e.g., images, text, numbers, and tables, we employ GPT-4 to automatically assess the answer.
-If the output is an image, we use gpt-4-vision-preview to compare the generated image to the label.
-Lastly, we calculate the total number of tokens consumed by the input prompt and output response, which denotes the solving efficiency of each method.
-Table 2: We demonstrate the accuracy and efficiency of all methods with gpt-3.5.
-Methods Exec@1(%) Pass@1(%) #Token
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 3: All samples are divided into three subsets based on its complexity. We compute Pass@1.
-Methods Sequential Parallel Complex Overall
-Vanilla 42.4% 29.0% 1.6% 28.5%
-ReAct 56.3% 48.6% 14.6% 44.1%
-Reflexion 67.8% 55.1% 48.1% 59.0%
-Multi-agent 63.2% 64.7% 35.5% 57.4%
-Data-Copilot 71.8% 70.0% 67.1% 70.2%
-4.2 Comparison Results
-As shown in Table 3, Data-Copilot surpasses all baselines in success rate (Pass@1), Exec@1, and Efficiency, achieving 11.2% improvements.
-Data-Copilot significantly reduces the risk of failure.
-Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%.
-In most cases, it merely needs to invoke pre-designed interface modules and generate corresponding arguments to address user requests, notably reducing the risk of errors.
-In contrast, baseline strategies must construct all code from scratch, making it easy to overlook certain details or make mistakes.
-Besides, the difference between Pass@1 and Exec@1 indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
-Data-Copilot reduces repetitive overhead for common requests.
-The consumed tokens represent computational overhead.
-Data-Copilot only consumes 561 tokens, just 30% of the baseline’s consumption.
-Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments.
-The primary token consumption lies in the input prompt which contains related interfaces descriptions.
-In contrast, baseline strategies may consume a lot of tokens since they need to output a complete solving code for each request.
-In the real world, numerous requests are similar or even repetitive.
-Data-Copilot can reduce repetitive overhead by abstracting common demands into interfaces.
-However, baseline methods have to repetitively generate code for each request.
-4.3 Analysis
-Data-Copilot exhibits a pronounced advantage on more challenging tasks.
-To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex.
-Complex workflows mean some intricate cases, mainly involving loops or nested workflows.
-As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement.
-However, for the more challenging parallel or complex tasks, our improvements are substantial, with enhancements of 5.3% and 19%, respectively.
-We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing.
-In contrast, our methods can effectively handle these tasks by invoking versatile interfaces and deploying workflow.
-These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
-We can deploy Data-Copilot locally using small LLMs with commendable performance.
-We analyze the performance of Data-Copilot with different LLMs.
-In Figure 4, we test GPT family and various open-source models.
-We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results.
-For instance, codellama-13B achieves a 50.8% Pass@1, almost similar to the best baseline in Table 3 (Reflexion with gpt-3.5-turbo).
-This indicates that Data-Copilot can be deployed locally at a relatively low cost while maintaining commendable performance.
-4.4 Human Evaluation
-To assess whether the solutions align with human preferences, we perform human evaluations across two dimensions: rationality and conciseness.
-Specifically, we collect the solutions of Data-Copilot and Reflexion.
-Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix B).
-We present results in Figure 4.
-It demonstrates that our solution significantly outperforms the competitor in terms of rationality.
-Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%.
-Besides, as for conciseness, Data-Copilot lags behind its competitor, with 27.8% wins against 30% losses.
-This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are universal, accommodating a wide array of functions, which makes our system less concise.
-The results of human evaluation are consistent with the previous assessment.
-Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
-4.5 Performance of other programming languages
-Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces.
+B.2 Human Evaluation on Test Benchmark
+We invite four more graduate students to manually evaluate our benchmark according to four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy.
+The belief guidance for human evaluation is as follows:
+  • Task Difficulty: The difficulty of the task, whether it requires multiple steps… Scoring Criteria: 5: very difficult,…, 1: easy.
+  • Request Rationality: Whether the request is reasonable, or if it is strange and does not align with human habits… Scoring Criteria: 5: Reasonable, aligns with humans..
+  • Expression Ambiguity: Whether the phrasing of the request is ambiguous, or if the entities involved are unclear… Scoring Criteria: 5: Clear without any ambiguity…
+  • Answer Accuracy: Whether the answer table is correct, detailed, and comprehensive, totally meeting user expectations… Scoring Criteria: 5: Data is totally correct … 3: Partial…
+Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
+B.3 Expanding to Other Programming Languages.
+In addition to the Python language, Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces.
 We tested three programming languages: Python, C++, and Matlab.
@@
 So ultimately, we opted for Python for Data-Copilot.
-Figure 3: Results of Data-Copilot using various LLMs as based model.
-LLM Pass@1 (%) #Token
-gpt-4 75.6 561.2
-gpt-3.5-turbo 70.2 559.3
-Llama-2-70B-chat 49.3 623.6
-codellama-13B 50.8 507.4
-text-davinci-003 47.5 495.9
-Llama-2-13B-chat 36.9 443.8
-Vicuna-13B-v1.5 33.2 502.8
-Figure 4: Win rate for Data-Copilot Vs. Reflexion on human preferences.
-5 Conclusion
-We propose a code-centric agent, Data-Copilot, to address a wide array of data-related tasks.
-Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests.
-It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
-Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
-Appendix
-Appendix A Experiments Details
-During the data exploration phase, we instruct LLM to generate Python code for each interface.
-Then we abstract interfaces from these requests.
-To merge the similar interfaces for each iteration, we adopt the following strategy:
-1. Firstly, we calculate the semantic similarity between the names of all existing interfaces and the newly designed interface, then sort them by similarity from highest to lowest.
-2.
-Then, we feed the top-10 existing interfaces into the LLM, and instruct the LLM to sequentially check whether they are functionally similar to the new interface, or can be merged using new parameters.
-3. Lastly, we output the merged interface. If the LLM deems no merging necessary, the new interface is retained.
-Besides, for the workflow deployment phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface.
-We set the temperature to 0.2 for more consistent results.
-Appendix B Human Evaluation Details
-We collect the solutions of Data-Copilot and Reflexion on 173 human-proposed requests.
-Subsequently, we involve experts to evaluate which of the two solving codes is more rational and concise.
-The criteria for expert are as follows:
-Rationality:
-1) Accurately identifies data of interest to the user.
-2) Processes the data logically.
-3) Output data in a format that meets human needs, with aesthetic and concise presentation.
-Conciseness:
-1) Only extracts and computes data relevant to the users interest.
-2) Code is clear, concise, and well-structured.
-3) The solution process is devoid of repetition and redundancy, ensuring a straightforward and intuitive approach.
-Experts are required to evaluate which of the two solutions is better based on these principles, and then we calculate the win rate of Data-Copilot against the Reflexion.
-Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
 Appendix C Limitations
@@
 We provide detailed prompts for our Data-Copilot and baselines.
-Specifically, for the data exploration, we provide detailed prompts in Section F.1 for four phases: self-request, interface design, and interface optimization.
+Specifically, for the data exploration, we provide detailed prompts in Section F.1 for four phases: self-exploration, interface design, and interface optimization.
 For the workflow deployment phase, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow.
-Additionally, we outline prompts for all baselines in Section F.3: Vanilla LLM, ReAct, Reflection, and Multi-Agent Collaboration strategies.
+Additionally, we outline prompts for all baselines in Section F.3: Direct-Code, ReAct, Reflection, and Multi-Agent Collaboration strategies.
 F.1 Prompts For Data Exploration
-Explore Data by Self-Request Phase
+Explore Data by self-exploration Phase
 ###Instruction: Given some data and its description, please mimic these seed requests and generate more requests.
@@
 F.3 Prompts For Baselines
-Vanilla LLM
+Direct-Code LLM
 ###Instruction: You are an artificial intelligence assistant.
@@
 Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The contribution of the article is clearly stated in Abstract and Section 1 (Introduction), and is fully consistent with the experiment (Section 3).
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss limitations in Appendix Section C.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This paper does not deal with the theoretical derivation part.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We have provided all the experimental details in the article (Section 4.1), prompt design (Appendix A, B, F.1 F.2 F.3).
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We have provided a code implementation for our system. The codebase has been open-sourced.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We have provided all the experimental details in the article (Section 4.1), prompt design (Appendix A, B, F.1 F.2 F.3).
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We provide details of the experiment, reporting the mean and standard deviation after multiple experiments in Table 2.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: In our experiments, we report the model of the API interfaces used and the tokens consumed for each experiment.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [N/A]
-  44. Justification: This paper does not involve any issues related to the Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: We propose an autonomous data analysis agent that does not generate any negative social impact.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [No]
-  54. Justification: Our method does not pose such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We utilized financial data provided by Tushare, which has been duly cited in the paper.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [No]
-  64. Justification: The paper does not release new assets
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [Yes]
-  69. Justification: We invited experts to annotate and evaluate the data, providing them with detailed instructions and guidelines.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: Our paper does not present these risks.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v8

@@
 Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily.
-Efficiently managing, processing, and displaying this data requires specialized expertise and is often tedious and repetitive.
+Efficiently managing, processing, and visualizing such data is labor-intensive and frequently necessitates specialized expertise.
 Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution.
-However, LLMs are not adept at handling complex numerical computations and table manipulations and are also constrained by a limited context budget.
-Based on this, we propose Data-Copilot, a data analysis agent that autonomously performs querying, processing, and visualization of massive data tailored to diverse human requests.
-The advancements are twofold: First, it is a code-centric agent that receives human requests and generates code as an intermediary to handle massive data, which is quite flexible for large-scale data processing tasks.
-Second, Data-Copilot involves a data exploration phase in advance, which explores how to design more universal and error-free interfaces for real-time response.
-Specifically, it actively explores data sources, discovers numerous common requests, and abstracts them into many universal interfaces (code modules) for daily invocation.
-When deployed in real-time requests, Data-Copilot only needs to invoke these pre-designed interfaces, transforming raw data into visualized outputs (e.g., charts, tables, and text) that best match the user’s intent.
-Compared to generating code from scratch, invoking these pre-designed and compiler-validated interfaces can significantly reduce errors during real-time requests.
-Additionally, interface workflows are more efficient and offer greater interpretability than code.
-We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news, demonstrating promising application prospects.
-Quantitative evaluations indicate that, compared to direct code generation and other agent strategies, our exploration-deployment strategy addresses human requests more accurately and efficiently, i.e., with higher success rates and lower token consumption.
-Figure 1: We present two strategies for using LLMs to automate data analysis.
-Upper: Directly using LLMs to read and process massive data is limited by context length, the LLM’s capabilities, and output format constraints.
+However, LLMs are not adept at handling complex numerical computations and table manipulations, and they are further constrained by a limited length context.
+To bridge this, we propose Data-Copilot, a data analysis agent that autonomously performs data querying, processing, and visualization tailored to diverse human requests.
+The advancements are twofold: First, it is a code-centric agent that leverages code as an intermediary to process and visualize massive data based on human requests, achieving automated large-scale data analysis.
+Second, Data-Copilot involves a data exploration phase in advance, which autonomously explores how to design universal and error-free interfaces from data, reducing the error rate in real-time responses.
+Specifically, It imitates common requests from data sources, abstracts them into universal interfaces (code modules), optimizes their functionality, and validates effectiveness.
+For real-time requests, Data-Copilot invokes these interfaces to address user intent.
+Compared to generating code from scratch, invoking these pre-designed and well-validated interfaces can significantly reduce errors during real-time requests.
+We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news.
+Quantitative evaluations indicate that our exploration-deployment strategy addresses human requests more accurate and efficiently, with good interpretability.
+Data-Copilot: Bridging Billions of Data and Humans with
+Autonomous Workflow
+Wenqi Zhang1, Yongliang Shen1, Zeqi Tan1, Guiyang Hou1, Weiming Lu1, Yueting Zhuang1 1College of Computer Science and Technology, Zhejiang University {zhangwenqi, luwm}@zju.edu.cn Project Page: https://github.com/zwq2018/Data-Copilot
+Figure 1: We compare two LLM strategies for automated data analysis.
+Upper: LLM’s capabilities, context length, and output format constraints limit the use of LLMs to process massive data.
 Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks.
-It actively explores how to design more universal and error-free interface modules, improving the success rate of real-time requests.
-Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow to fulfill human intent.
+It explores how to design more universal and error-free interface modules, improving the success rate of real-time requests.
+Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow for human intent.
 1 Introduction
@@
 Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption.
-However, these data-related tasks often require specialized knowledge and tedious manual labor.
-Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022), especially when combined with techniques such (Wei et al., 2022b; Kojima et al., 2022) has demonstrated their capability to handle complex tasks.
-Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow that performs data analysis and outputs visualization results that best match user expectations?
-An intuitive solution is to directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023), just as we do with text.
-However, several challenges must be considered: (1) Due to the input context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text.
+However, these data-related tasks often require tedious manual labor and specialized knowledge.
+Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022) and techniques (Wei et al., 2022b; Kojima et al., 2022) have demonstrated the capability to handle complex tasks.
+Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow performing data analysis and visualization that best matches user expectations?
+An intuitive solution is to treat data as a special text, i.e., directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023).
+However, as shown in Figure 1, several challenges must be considered: (1) Due to context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text.
 Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources.
-(2) Data processing is very complex, involving many tedious numerical calculations and intricate table manipulations.
+(2) Data processing is complex, involving many tedious numerical calculations and intricate table manipulations.
 LLMs are not adept at performing these tasks.
-(3) An LLM can only output text, whereas we usually prefer charts or tables when analyzing data.
-These challenges significantly constrain the application of LLMs in data-related tasks.
-Recently, many works have explored alternative solutions.
-LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration.
-Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024) and DB-GPT (Xue et al., 2023) apply LLMs to data science domain like Text2SQL.
-Besides, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b).
-Data Interpreter (Hong et al., 2024) proposes a code-based agent for data science through the plan-code-verify paradigm.
+(3) Data analysis typically requires visualizing the results of data processing, whereas LLMs are limited to generating text output.
+These challenges constrain the application of LLMs in data-related tasks.
+Recently, many agent-based designs have explored alternative solutions (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b; Nejjar et al., 2023; Li et al., 2024b).
+LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on exploring insight from data.
+Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024b), DB-GPT (Xue et al., 2023) and TAG (Biswal et al., 2024) apply LLMs to data science domain like Text2SQL.
+Data Interpreter (Hong et al., 2024) proposes a plan-code-verify paradigm for automating machine learning tasks.
 These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
 Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent.
-Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and output multi-modal results (e.g., chart, table, text) for humans, reducing tedious human labor and eliminating the need for direct data access by the LLM.
-However, creating a reliable code agent for various data analysis tasks is also not an easy feat.
-Firstly, LLM-generated code often contains formatting errors, logical inconsistencies, and code hallucinations, such as fabricating non-existent functions, which may prevent it from being deployed in real-world environments.
-Secondly, most code-based agent generates code for solving user requests, without considering the relations between different requests and only addressing each issue independently.
-Based on this observation, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis.
-First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches the user’s intent.
+Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and visualize its results (e.g., chart, table, text) for humans.
+However, creating a code agent that can be used in real-world data analysis tasks is far from an easy feat.
+① LLMs struggle to generate high-quality, error-free code in a single attempt, often containing format errors, logical inconsistencies, or fabricating non-existent functions.
+② Although the inference speed of LLMs has significantly improved, generating lengthy code still consumes a considerable amount of time and tokens.
+These two challenges—high code error rate and inefficient inference—must be addressed urgently.
+To address this, we observe most human requests are either similar or inherently related.
+By abstracting common demands into interfaces and validating their functionality in advance, we can significantly improve both the success rate and efficiency of real-time deployment.
+Therefore, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis.
+First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches user’s intent.
 The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks.
-Besides, Data-Copilot also incorporates a data exploration phase.
-During this phase, it actively explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance.
-In a real-world deployment, Data-Copilot flexibly invokes pre-design interface modules to handle user requests, creating an interface workflow that transforms raw data into desired visualization results.
-This design can bring three advantages.
-Firstly, this exploratory process allows Data-Copilot to analyze connections between data, design general interfaces for similar requests, and pre-validate their correctness, thereby reducing errors in real-time responses.
-Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving generation efficiency.
-Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
-As shown in Figure 1, Data-Copilot contains three key steps for data exploration:
-  1. 1.
-    Explore data and synthesize requests: Data-Copilot discovers potential requests and a broader range of human need scenarios from data.
-It involves a "self-exploration" process to generate massive synthesized requests based on all data schemas and seed requests.
-  2. 2.
-    Interface design and test: Data-Copilot generates modular code for request solving. Each interface is a functional module, and test cases are generated to verify the designed interface.
-  3. 3.
-    Interface optimization: To improve the versatility of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces based on the compiler’s feedback.
-After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks.
+Besides, Data-Copilot also incorporates a data exploration and interface design phase.
+It autonomously explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance.
+In real-world deployment, Data-Copilot flexibly invokes pre-design interface modules for most requests, deploying a well-verified interface workflow for data processing and visualization.
+Data-Copilot brings three advantages.
+Firstly, this exploratory process allows Data-Copilot to analyze and summarize the inherent connections between human requests, design general interfaces for similar requests, and pre-validate their correctness, reducing errors in real-time responses.
+Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving inference efficiency.
+Lastly, compared to lengthy code, these interfaces provide greater interpretability, since it is easier for human reading and interaction.
+To achieve this, we contains three steps (Figure 1) when self exploration:
+Explore data and Synthesize requests: Data-Copilot discovers potential requests and a broader range of human need from data.
+It involves a "self-exploration" process to generate massive requests based on all data schemas and seed requests.
+Interface Design and Test: It designs modular interface from synthesized requests, with test cases automatically generated for verification.
+Interface Optimization: To improve versatility, it merges similar interfaces and also revises erroneous interfaces using compiler’s feedback.
+After exploration-design-optimization, Data-Copilot designs many general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks.
 When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests.
 For different requests, Data-Copilot can flexibly deploy various invocation structures, such as step-by-step serial workflows, parallel, or loop.
-It can even output a hybrid form of interface workflows and raw code for some "unfamiliar" requests.
+It can even output a hybrid form of interface workflows and raw code for these "unfamiliar" requests.
 Our contributions are threefold:
@@
 Data Exploration: First, it performs a self-exploration process to uncover potential human requests from data sources.
-Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces one by one.
+Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces.
 This exploration-design-optimization process is operated in advance.
-Workflow Deployment: Upon receiving real-time requests, Data-Copilot can quickly invoke existing interfaces and deploy a workflow for familiar requests, or flexibly combine interfaces and raw code for "uncovered" requests.
-  1. 1.
-    We propose a code-centric agent, Data-Copilot, for massive data analysis tasks.
-It leverages LLM and its code generation capabilities to automate data querying, processing, and visualization tailored to human requests, reducing tedious human labor.
-  2. 2.
+Workflow Deployment: Upon receiving real-time requests, Data-Copilot invokes existing interfaces and deploys a workflow for familiar requests, or flexibly combines interfaces and raw code for "uncovered" requests.
+  • We propose a code-centric agent, Data-Copilot, for automated data analysis and visualization.
+It leverages LLM’s code generation abilities for data querying, processing, and visualization, reducing tedious human labor.
   • We decouple code generation into two phases: data exploration and workflow deployment.
-In the exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data.
-These interfaces enhance the success rate and efficiency of real-time responses.
+In exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data.
 In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests.
-  3. 3.
-    We open-source Data-Copilot for diverse Chinese financial data analysis, including stocks, funds, and live news.
-Quantitative evaluations indicate that our agent outperforms other agent strategies, with higher success rates and lower token consumption.
-Besides, interface workflows are more convenient for human inspection and interaction, offering stronger interpretability.
+This design enhances the success rate and efficiency of real-time responses.
+  • We open-source Data-Copilot for Chinese financial data analysis, including stocks, funds, and live news.
+Quantitative evaluations indicate our agent outperforms other strategies, with higher success rates and lower token consumption.
+Besides, interface workflows are more convenient for human inspection and interaction, offering interpretability.
 2 Data-Copilot
-Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent.
-Data-Copilot operates in two phases: Data Exploration and Workflow Deployment.
+Data-Copilot is a code-centric agent capable of performing data analysis and visualization based on human instructions.
+It operates in two phases: Data Exploration and Workflow Deployment.
 In the first phase (Section 2.1), Data-Copilot designs a self-exploration process to discover numerous potential needs based on data schemas (Step 1: exploration).
-Then based on these synthesized requests, it abstracts universal interface modules for each request (Step 2: design).
+Then based on these synthesized requests, it abstracts many universal interface modules (Step 2: design).
 After that, it optimizes similar interfaces and tests their correctness, ensuring each interface is correct and universal (Step 3: testing and optimization).
-The whole exploration-design-optimization process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
+The whole process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
 When faced with real-time requests (Section 2.2), Data-Copilot can flexibly invoke pre-designed interfaces or directly generate raw code based for the user’s request.
 In most cases, existing interfaces can cover the majority of real-world requests, significantly enhancing both success rate and response speed.
-We provide a detailed prompt for two phases in Section F.1 and F.2.
+We provide a detailed prompt for two phases in Section E.3 and E.4.
 2.1 Data Exploration
 Let’s review how human data analysts operate.
-Initially, they need to observe what data is available, understand the data formats, and determine how to use them.
-Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability.
+Initially, they need to observe the available data, understand the data formats, and learn how to access them.
+Subsequently, humans often design generic modules to simplify the code logic and test these modules for usability.
 Similar to this, Data-Copilot also autonomously explores data and derives insights from vast data sources, including the relations between the data, and the potential requests associated with the data.
@@
 This process of exploring data, identifying common requests, designing general interfaces, and testing and optimizing their performance is conducted in advance on its own.
-Self-Exploration Request
+Self-Exploration Request.
 To explore data and mine insights, we design a self-exploration process.
@@
 Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example.
-Then we feed the parsing files and a few seed requests into the LLMs, and prompt LLMs to generate more diverse requests based on these data, covering different demand scenarios as comprehensively as possible.
-The brief prompt is shown as follows:
-###Instruction: Given some data and its description, please mimic these seed requests and generate more requests.
-The requests you generate should be as diverse as possible, covering more data types and common needs.
-###Seed Request: {request1, request2,...}
-###Parsing file for GDP_Data:{
-Description: This data records China’s annual and quarterly GDP...,
-Access Method: pro.cn-gdp(start-time, end-time, frequency,...),
-Output Schema: Return 9 columns, including quarter, gdp, gdp-yoy...,
-Usage:{
-Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’,..),
-First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
-###Parsing file for Stock_data:{...}, ....
-How to ensure the quality of the generated requests?
+Then we feed the parsing files and a few seed requests into the LLMs, and prompt LLMs to synthesize more diverse requests based on these data.
+The brief example is shown in Section E.2.
+The quality of the generated requests.
 A common issue is that the LLM often proposes a request about the data that doesn’t exist.
-To address this, we design a backward verification process to check the rationality of these requests.
+To address this, we design a backward verification strategy to check these synthesized requests.
 Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests.
-Upon manually checking some requests, we found that the generated requests generally met our expectations.
-Interface Design
+Besides, when synthesizing, we also use keywords to control the topics of synthesized requests, ensuring they closely align with real-world distribution.
+Upon manual evaluation of the synthesized requests, we found that the generated requests generally met our expectations.
+We discuss the quality of synthesized requests in detail in Tables 1 and B2.
+Interface Design.
 After generating massive requests through self-exploration, Data-Copilot abstracts many universal interfaces from these requests.
@@
 It performs specific tasks such as data retrieval, computation, and visualization.
-In our paper, each interface is designed, tested, and optimized iteratively by Data-Copilot.
-First, starting from the initial request, we iteratively feed self-exploration requests and related data parsing files into LLMs, prompting the LLM to design complete code modules for request solving.
-Each code module corresponds to an interface and is stored in the interface library.
+Each interface is designed, tested, and optimized iteratively by Data-Copilot.
+First, starting from the initial request, we iteratively feed synthesized requests from the self-exploration stage and related data parsing files into LLMs, prompting the LLM to design complete code modules for request solving.
+Each code module is defined as an interface and stored in the interface library.
 During each iteration, LLMs are instructed to prioritize utilizing the existing interfaces within the library.
 If the available interfaces are insufficient for request solving, the LLMs design a new interface.
-As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...).
-As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
-Interface Testing
-After designing a new interface for a request (request-i $\rightarrow$ interface-i), Data-Copilot autonomously tests its correctness based on compiler feedback.
-First, Data-Copilot uses the request-i as the seed to generate massive similar requests as test cases.
-Then it tests the new interface-i one by one.
-E.g., in Fig.
-2, for newly designed interface-i: Query-GDP(), Data-Copilot mimics request-1 to generate many similar requests to test the Query-GDP(): "I want to see USA’s GDP over the past 10 years", "I want to see China’s GDP for the last year".
-If the interface passes all test cases, it is retained; otherwise, Data-Copilot utilizes compiler error feedback for self-reflection, correcting the erroneous code snippets until it successfully passes the tests.
-Interface Optimization
-To further improve the generality of interfaces, Data-Copilot also optimizes similar interfaces.
-Similar to how human software developers, each time a new interface is designed, this optimization process is triggered: evaluating whether the newly designed interface can be merged with previous ones.
+As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP() and Plot-Line().
+As for the second request: Compare the CPI of China and the USA over.., since the previous two could not solve this request, a new interface, Query-CPI(), is designed by LLMs.
+[…]
+Interface Optimization.
+To optimize the generality of designed interfaces, Data-Copilot also merges similar interfaces.
+Similar to human developers, each time a new interface is designed, this optimization process is triggered: evaluating whether the newly designed interface can be merged with previous ones.
   • Retrieve similar interfaces.
-After a new interface is designed, we first retrieve the top N similar interfaces from the existing interface library.
-Specifically, we use gte-Qwen1.5-7B-instruct (Li et al., 2023d) to obtain embeddings of each interface code and then calculate their similarity.
+After a new interface is designed, we retrieve Top-N interfaces from existing interface library.
+Specifically, we use gte-Qwen1.5-7B-instruct (Li et al., 2023d) to obtain embeddings of each interface code and then calculate their similarity, identifying top N similar interfaces.
   • Decide whether to merge.
-Then LLMs are prompted to evaluate the functionality, parameters, and logic of the new interface and the top N similar interfaces, autonomously deciding whether to merge the new interface with existing ones.
-If the LLM deems no merging necessary, the new interface is retained.
-As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...).
-This optimization process makes each interface more general and unique, reducing interface redundancy.
+LLMs are prompted to compare new interface with Top-N retrieved interfaces in terms of functionality, parameters, and processing logic, autonomously deciding whether to merge it with the existing ones.
+If LLM deems no merging necessary, new interface is retained.
+As shown in Figure 2, two interfaces Query-GDP() and Query-CPI() are merged into Query-GDP-CPI().
+This process makes each interface more general and unique.
   • Test the optimized interface.
 When two interfaces are merged, Data-Copilot also needs to test the newly merged interface.
-The test cases from the two original interfaces are used to validate the merged interface.
-The output of the new interface must be consistent with the two original interfaces in both format and content.
-For each request, Data-Copilot alternates the above three steps: interface design, testing, and optimization until all the requests can be covered by these interfaces.
-As shown in Figure G3, Data-Copilot designs many interfaces for different task and operation types, e.g., data acquisition, prediction, visualization, and DataFrame manipulation.
-We provide some detailed cases in Section G.1.
+Test cases from two original interfaces are used to validate the merged interface.
+The output of the new interface must be consistent with two original interfaces in both format and content.
+In this phase, Data-Copilot alternates the above three steps: interface design, testing, and optimization until all the requests can be covered by these interfaces.
+As shown in Figure F3, Data-Copilot designs many interfaces for different task and operation types, e.g., data acquisition, prediction, visualization, and DataFrame manipulation.
+We provide detailed algorithm processes (Section E.1) and examples in Section F.1.
 2.2 Workflow Deployment
-In the first phase, Data-Copilot designs many generic interfaces with different functionalities in advance.
-They are versatile and error-free, capable of handling the majority of human requests in real-world scenarios.
-As shown in Figure 2, Data-Copilot only needs to invoke the relevant interfaces for human requests, which is more accurate and efficient (interface workflow).
-Furthermore, for some long-tail or new requests that cannot be covered by the existing interfaces, Data-Copilot can flexibly generate raw code to handle them (interface-code hybrid strategy) and document these requests for updating the interface library in the future.
-Task Selection and Interfaces Retrieval
+As Figure 2 shows, it accurately and efficiently handles requests by invoking relevant interfaces (interface workflow).
+For long-tail or uncovered requests, it flexibly generates raw code (interface-code hybrid strategy) while documenting these for future interface library updates.
+Interfaces Retrieval.
 Considering the significant differences between the interfaces involved in different requests, it is unnecessary to load all interfaces every time.
@@
 Upon receiving a real-time request, Data-Copilot first determines the appropriate task types and required operation types and then loads the interfaces associated with these types for subsequent workflow planning.
-This can significantly reduce the number of interfaces in the prompt.
-Planning Workflow with Different Structures
-After reading the descriptions of all retrieved interfaces, Data-Copilot plans an interface workflow to process the user request.
-A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex loop structures.
-Therefore, Data-Copilot needs to autonomously determine which interfaces to invoke, in what sequence, and which parameters to populate.
-The interface invocation workflow is output in JSON format: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}.
-The prompts and cases are shown in Sections F.2 and G.2.
-As shown in Figure 2, Data-Copilot designs interface workflows with different structures, e.g., sequential, parallel, or loop.
-For instance, the first request in Figure 2: "Compare the return of CSI-300, GEM and CSI-1000 this year", Data-Copilot plans a parallel workflow with 5 steps: Data Acquisition(), Calcuate Return() in parallel for calculating three indices, followed by Visualization().
-In the second case, it deploys a loop workflow by calling the Loop().
-Interface-Code Hybrid Generation
+This can reduce the number of interfaces in the prompt.
+Interface Invocation Workflow.
+After reading interface descriptions, Data-Copilot plans workflows—multiple interfaces in specific order forming chain, parallel, or loop structures.
+It determines which interfaces to invoke, their sequence, and parameters, outputting in JSON format.
+Prompts and cases appear in Sections E.4 and F.2.
+As Figure 2 shows, Data-Copilot designs sequential, parallel, or loop interface workflows.
+For request "Compare the return of CSI-300, GEM and CSI-1000 this year", it plans a parallel workflow: Data Acquisition(), Calculate Return() for three indices simultaneously, then Visualization().
+The second case implements a loop workflow using Loop().
+Interface-Code Hybrid Generation.
 In the real world, it is inevitable to encounter "uncovered" requests that cannot be addressed by existing interfaces.
 As a remedy, Data-Copilot adopts an interface-code hybrid generation strategy.
-Specifically, Data-Copilot prioritizes invoking existing interfaces to resolve user requests.
+Specifically, it prioritizes invoking existing interfaces to resolve user requests.
 It can solve most common user requests.
-However, if the deployed workflow continues to fail, or if Data-Copilot proactively determines that the current request cannot be resolved by existing interfaces, our system is allowed to generate raw code directly or generate a combination of raw code and interfaces for request solving.
+However, if the deployed workflow continues to fail, or if Data-Copilot proactively determines that current request cannot be resolved by existing interfaces, it would generate raw code directly or generate a combination of raw code and interfaces.
 This design endows Data-Copilot with the flexibility to handle diverse requests.
 Besides, each time Data-Copilot encounters such "uncovered" requests, it also records them in the document.
-After accumulating a certain number of new requests, we re-initiated the Data Exploration stage (Section 2.1) to develop new interfaces for these newly emerged requests.
-It empowers Data-Copilot to continuously update its interface library throughout human interactions.
+After accumulating sufficient new requests, we re-initiated the Data Exploration and interface design stage (Section 2.1), i.e., in real-world interactions, we periodically develop new interfaces for emerging demands, continuously updating Data-Copilot’s interface library.
+Name Source #Cases Form Four Types Ratio Three Complexity Levels Ratio
+Seed Set Human-proposed 173 Query 8.0: 4.0: 4.0: 1 2.0: 1.5: 1
+Set for Interface Design Self-Exploration 3000 Query 8.7: 3.3: 4.2: 1 1.9: 1.5: 1
+Test Set Self-Exploration & Human 547 Query, Label 8.6: 3.3: 4.3: 1 1.8: 1.4: 1
 Table 1: Statistics on human-proposed requests and self-exploration dataset.
-Source #Cases Form Usage Ratio of 4 Tasks Ratio of 3 Complexities
-Human-proposed 173 Query Seed Set 8: 4: 4: 1 2.0: 1.5: 1
-Self-Exploration 3000 Query Interface Design 8.7: 3.3: 4.2: 1 1.9: 1.5: 1
-Self-Exploration & Human 547 Query, Label Test Set 8.6: 3.3: 4.3: 1 1.8: 1.4: 1
-3 Qualitative Evaluation
+We report the number of each task type and complexity level.
+The results indicate that our synthesized requests closely align with human distribution.
+3 Dataset Synthesis
 3.1 Environment and Data Sources
-Data-Copilot is evaluating on Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data.
-Similar to many real-world financial companies, Data-Copilot utilizes data interfaces provided by Tushare [^1] https://tushare.pro/ to access vast amounts of financial data, including time-series data spanning over 20 years for more than 4,000 stocks, funds, and futures.
-In the first phase, we use a strong LLM (e.g., gpt-4-turbo) for data exploration and a lightweight LLM (e.g., gpt-3.5-turbo) for workflow deployment.
+Data-Copilot is developed on Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data.
+Similar to many works, Data-Copilot utilizes data interfaces provided by Tushare [^1] https://tushare.pro/ to access vast amounts of financial data, including time-series data spanning over 20 years for more than 4,000 stocks, funds, and futures.
+In the first phase, we use a strong LLM (e.g., gpt-4o) for data exploration and a lightweight LLM (e.g., gpt-3.5-turbo) for workflow deployment.
 3.2 The Creation of Dataset
-Human-proposed Request as Seed.
+Human-proposed Request as Seed Set.
 We invite 10 students in economics to submit 50 requests each.
-These requests cover a wide range of common needs, including stock tasks, fund tasks, and corporate tasks, and also involve different difficulty levels.
-We treat these 500 human-proposed requests as a distribution of real-world scenarios, where the ratio of four task types is Stock: Fund: Corporation: Other = 8: 4: 4: 1.
-Then we filter out highly similar requests by computing their sentence embedding.
-Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
+These requests cover a wide range of common needs, including stock, fund tasks with different complexity levels.
+These human-proposed requests can represent a distribution of real-world scenarios, where the ratio of four tasks is: Stock(8), Fund(4), Corporation(4) and Others(1).
+Then we filter out highly similar requests.
+Lastly, we retain 173 high-quality requests, which are used as seeds for Data Exploration phase.
 Self-Exploration Request Set.
 As mentioned in Section 2.1, we employ self-exploration to expand the request set.
-Specifically, we feed 173 human-proposed seed requests along with all data descriptions and schema into GPT-4 for data exploration and request synthesis.
-When self-exploration, we use keywords to control the distribution of the synthesized requests, ensuring that they can closely align with real-world scenarios.
-In this phase, we generate 6480 requests.
-After that, we filter out highly similar requests and retain 3547 requests.
-Then we adopted a stratified sampling strategy to sample 547 instances as the test set from each task type.
+We first feed 173 human-proposed seed requests along with all data descriptions and schema into GPT-4 for data exploration and request synthesis.
+We use keywords to control the distribution of the synthesized requests, ensuring close alignment with the real world.
+When exploration, it generates 6480 requests.
+Then we filter out highly similar requests and retain 3547 requests.
+Then we adopt a stratified sampling strategy to sample 547 instances as test set from each type.
 The remaining 3000 requests are used for interface design.
-We analyze the number of subsets categorized by four task types (stock, fund, corporation, and others) and three complexity levels (single-entity, multi-entities, and multi-entities with complex relations).
-The results indicate that our synthesized requests closely align with human distribution.
-The detailed statistics of three datasets are shown in Tables 1, B1 and B1.
-Table 2: Human evaluation on the human-proposed seed set and self-exploration requests.
-Number Task Difficulty Request Rationality Expression Ambiguity Answer Accuracy
-Seed Set 173 3.5 4.3 4.4 -
-Test Set 547 3.9 4.2 4.0 4.8
+The distributions of four task types (stock, fund, corporation, and others) and three complexity levels (single-entity, multi-entities, and multi-entities with complex relations) are shown in Table 1.
+It shows that synthesized requests closely align with humans.
+The detailed statistics are shown in Tables 1, B1 and B1.
+Model GPT-4o GPT-4-turbo GPT-3.5-turbo Llama3-70B-Instruct Llama2-70B-Instruct
+Direct-Code 73.6 70.0 28.5 52.2 29.6
+Ours 77.9 $\uparrow$4.3 74.6 $\uparrow$4.6 70.2 $\uparrow$41.7 70.9 $\uparrow$18.7 49.3 $\uparrow$19.7
+Model Codellama-13B Vicuna-13B-v1.5 DeepSeek Coder-V2 Qwen2.5-Coder-32B Qwen2.5-Coder-14B
+Direct-Code 21.0 17.3 63.8 66.5 51.7
+Ours 50.8 $\uparrow$29.2 33.2 $\uparrow$15.9 75.3 $\uparrow$11.5 77.2 $\uparrow$10.7 72.3 $\uparrow$20.6
+Table 2: Accuracy of ours and direct code generation (Direct-Code) using different LLMs for workflow deployment.
 Annotate Answer Table for Test Set.
-Human annotators are instructed to annotate data-table as answers for 547 test requests, which is used for benchmarking our system’s output.
-Specifically, annotators manually retrieve and process the corresponding data based on the testing request, and record the final data table before chart plotting.
-These data tables are used as answer annotations.
-For example, for the request: "I want to see China’s GDP trend over the past 5 years", the answer data-table is "[2023: 17.8 trillion, 2022: 17.9 trillion, 2021: 17.8 trillion, 2020: 14.6 trillion, 2019: 14.3 trillion]".
-The Quality of Self-exploration Requests.
-To assess the quality of self-exploration requests, we invited four additional graduate students to manually evaluate our test set (546 requests) and human-proposed requests (173 seed set) on four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy.
-We provide detailed guidance for each criterion in Section B.2.
-The results of two sets are shown in Table 2.
-We observed that the synthesized requests exhibit a comparable quality to human-proposed requests, with slightly higher difficulty, ambiguity, and similar rationality.
-It ensures that our test set can reflect most of the real-world demands.
-Besides, the evaluators gave a high score of 4.8 on the label (answer table) of our test set, which also ensures the accuracy of our dataset.
-3.3 Qualitative Results
-The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request.
-The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces.
-For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
-Data-Copilot designs versatile interfaces through data exploration.
-After exploring the Chinese financial data and processing 3000 self-exploration requests, Data-Copilot designed a total of 73 interfaces, comprising 5 functionalities and 4 tasks: data acquisition, index calculation, table manipulation, visualization, and general processing logic.
-These tasks include stocks, funds, corporations, and others.
-We present some representative interfaces and the statistics of the interfaces in Figure G3.
-Most interfaces are highly versatile with multiple functions.
-For example, get-stock-prices-data() interface can retrieve data on a daily, weekly, or monthly basis.
-These versatile interfaces endow the Data-Copilot with the capability to address complex requests.
-Data-Copilot can flexibly deploy interfaces or raw code for diverse demands.
-As mentioned in  Section 2.2, for some "new" requests, Data-Copilot is capable of addressing them by combining both raw code and existing interfaces.
-As shown in Figure 2 (3rd example), the user requests a radar chart, which was not directly supported by the available interface.
-Data-Copilot first leverages the existing interface to retrieve and process the data.
-Then based on the dataframe produced by the last interface, it generates raw code to render a radar chart.
-This design grants Data-Copilot greater flexibility to handle the constantly emerging new requests.
-Interface Workflow is human-friendly, offering higher interpretability.
-Upon executing the workflow via the compiler, Data-Copilot produces the visualized results.
-Also, it provides a comprehensive summary of the entire processing process based on the workflow.
-As shown in Figures 2, G2, G4, G6, G5 and G7, Data-Copilot’s outputs (workflow, visualized results and summary) are highly intuitive and clear.
-It encapsulates complex code processing logic into a step-by-step interface invocation workflow, which is easier for humans to understand and inspect.
-Table 3: Pass@1 of Data-Copilot and direct code generation (Raw Code) when using different LLMs for workflow deployment: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Llama-3-70B-chat, Llama-2-70B-chat, Llama-2-13B-chat, Codellama-13B, Vicuna-13B-v1.5).
-GPT-4o GPT-4 GPT-3.5 Llama3-70B Llama2-70B Llama2-13B Codellama13B Vicuna
-Raw Code 73.6 70.0 28.5 52.2 29.6 17.1 21.0 17.3
-Ours 77.9 $\uparrow$4.3 74.6 $\uparrow$4.6 70.2 $\uparrow$41.7 70.9 $\uparrow$18.7 49.3 $\uparrow$19.7 36.9 $\uparrow$19.8 50.8 $\uparrow$29.2 33.2$\uparrow$15
+Human annotators are instructed to annotate data tables as answers for 547 test requests.
+Specifically, annotators manually retrieve and process the corresponding data based on testing request, and record the final data table before chart plotting.
+For example, request: "I want to see China’s GDP over past 5 years", the labeled data-table is [2023: 17.8 trillion, 2022: 17.9 trillion, 2021: 17.8 trillion, 2020: 14.6 trillion, 2019: 14.3 trillion].
+The quality of self-exploration requests. We manually evaluate the quality of testset. As shown in Section B.3, these synthesized requests show a comparable quality to human-proposed requests.
+3.3 Qualitative Evaluation
+As shown in Figure F9, Data-Copilot constructs the invocation workflow step-by-step (each step corresponds to one or more interfaces) and final results (bar, chart, text) for input request.
+Data-Copilot designs versatile interfaces via data exploration.
+After analyzing Chinese financial data and 3,000 self-exploration requests, Data-Copilot created 73 interfaces across five functionalities (data acquisition, index calculation, table manipulation, visualization, general processing).
+Figure F3 shows key interfaces.
+Data-Copilot adapts to new requests by combining interfaces and raw code.
+As in  Section 2.2, for unsupported requests, Data-Copilot integrates existing interfaces with raw code.
+In  Figure 2 (3rd example), it retrieves data via interface, then generates radar chart code, enhancing flexibility for evolving demands.
+Interface Workflow enhances interpretability.
+After execution, Data-Copilot generates visuals and workflow summaries.
+As shown in Figures 2, F2, F4, F6, F5 and F7, outputs are intuitive.
+Structuring complex code into step-by-step interface calls improves clarity and inspection ease.
 4 Quantitative Evaluation
 4.1 Experiments Settings
-Baselines and Experiments Details We first compare Data-Copilot with direct code generation (denoted as Direct-Code) using different LLMs.
-Besides, we also compare Data-Copilot with many prompting and agent strategies, including: 1.
-ReAct (Yao et al., 2022): LLMs use neural language (thought) and code to resolve the user request step-by-step.
-2.
-Reflexion (Shinn et al., 2023): LLM can reflect on the feedback from the compiler at each step, progressively fulfilling user requests.
-For the Reflexion strategy, we limit it to a maximum of two rounds of reflection.
-3.
-Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion.
-For a fair comparison, Data-Copilot employs the same LLM as all baselines for evaluation when invoking pre-designed interfaces.
-The detailed prompts of all baselines are shown in Section F.3.
+Baselines and Experiments Details We compare Data-Copilot with direct code generation (Direct-Code), ReAct, Reflexion, Multi-agent methods. Detailed prompts are in Sections B.1 and E.5.
 Evaluation Methods.
-All methods are evaluated on the 547 testing instances.
-We focus on the accuracy of the final results and the efficiency of the response.
-Considering that the output of most requests is images or charts, we evaluate from two perspectives: whether the data used for plotting is correct and whether the drawn image meets human requirements.
-We use GPT-4-turbo and GPT-4o to conduct automatic evaluations.
-(1) Data-table Evaluation: We first obtain the dataframe before visualization, then use GPT-4-turbo to compare this data table with the answer table (labeled by humans in Section 3.2).
-(2) Image Evaluation: To evaluate whether the final image meets human requests, we feed the human request, human-annotated answer table, and the output image into GPT-4o for evaluation.
-We instruct GPT-4o to check whether visual elements in the image meet the human requests, including numerical points, lines, axes, image aesthetics, and style.
-Finally, the results of data and image evaluation are combined as Pass@1.
-Lastly, we calculate the total number of tokens consumed by each method (denoted as #Token), which represents the response efficiency.
-Table 4: We demonstrate the accuracy and efficiency of all methods with gpt-3.5-turbo
-Methods Pass@1(%) #Token
+We evaluate all methods on 547 test cases, focusing on three aspects: data-table accuracy, image quality, and inference efficiency.
+Evaluations are conducted using GPT-4o: Data-Table Evaluation: GPT-4o/turbo compares the generated dataframe with a human-labeled answer table (Section 3.2).
+Image Evaluation: To evaluate whether the final image meets human requests, we feed the human request, human-annotated answer table, and the generated image into GPT-4o.
+GPT-4o is instructed to check image’s visual elements based on human request, including numerical points, lines, axes, image aesthetics, and style, i.e., we design a checklist containing 5 categories with 10 sub-dimensions for GPT-4o scoring.
+The results of data-table and image evaluation are combined as Accuracy.
+Efficiency: Measured by the total token consumption (#Token) per method, representing solving efficiency.
+The evaluation details and comparison between manual evaluation and model evaluation are shown in Section B.4.
+Methods Accuracy(%) #Token
 Direct-Code 28.5 $\pm$2.1 823
-ReAct 44.1 $\pm$3.4 1515
-Reflexion 59.0 $\pm$3.9 2463
-Multi-Agent 57.4 $\pm$1.8 2835
+ReAct (Yao et al., 2022) 44.1 $\pm$3.4 1515
+Reflexion (Shinn et al., 2023) 59.0 $\pm$3.9 2463
+Multi-Agent (Hong et al., 2023) 57.4 $\pm$1.8 2835
 Data-Copilot 70.2 $\pm$2.5 561.2
@@
 Data-Copilot + Reflexion 71.8 $\pm$2.2 978
-Table 5: All samples are divided into 3 complexity levels. We compute Pass@1.
-Methods Single Multiple Complex Overall
-Direct-Code 42.4% 29.0% 1.6% 28.5%
-ReAct 56.3% 48.6% 14.6% 44.1%
-Reflexion 67.8% 55.1% 48.1% 59.0%
-Multi-agent 63.2% 64.7% 35.5% 57.4%
-Data-Copilot 71.8% 70.0% 67.1% 70.2%
+Table 3: Accuracy and efficiency on gpt-3.5-turbo.
+Methods Single Multiple Complex Rel. Overall
+Direct-Code 42.4 29.0 1.6 28.5
+ReAct 56.3 48.6 14.6 44.1
+Reflexion 67.8 55.1 48.1 59.0
+Multi-agent 63.2 64.7 35.5 57.4
+Data-Copilot 71.8 $\uparrow$4 70.0 $\uparrow$5.3 67.1 $\uparrow$19 70.2 $\uparrow$11.2
+Table 4: Accuracy for three complexity levels samples.
 4.2 Comparison Results
-Data-Copilot Significantly Reduces the Risk of Failure When Real-time Deployment.
-As demonstrated in Table 3, Compared to directly generating code, Data-Copilot achieves significant improvements across all LLMs, with particularly strong gains of +41.7, +18.7, and +29.2 on GPT-3.5-turbo, Llama3-70B, and Codellama-13B, respectively.
-By decoupling direct code generation into two stages: interface design in advance and real-time workflow deployment, Data-Copilot can elevate many smaller LLMs (e.g., Lllama3-70b, GPT-3.5) to GPT-4 level during real-time deployment.
-Through careful comparison, we observe that these pre-designed interfaces significantly reduce the occurrence of errors during real-time responses, as they have already been optimized and validated during the design phase.
-In contrast, directly generating code from scratch makes it easy to overlook certain details or make mistakes.
-Besides, we also found that even with a powerful LLM like GPT-4, Data-Copilot also achieves an improvement of over +4%, validating its generalizability.
+Data-Copilot Significantly Reduces Deployment Failure Risk.
+As shown in Table 2, compared to direct code generation, Data-Copilot achieves significant improvements: +41.7 (GPT-3.5-turbo), +18.7 (Llama3-70B), and +29.2 (Codellama-13B).
+By decoupling into interface design and workflow deployment stages, smaller LLMs (Llama3-70B, GPT-3.5) perform at GPT-4 level during deployment.
+Even GPT-4 improves by +4%, confirming generalizability.
+Pre-designed interfaces reduce errors through optimization and validation during design, whereas direct code generation often overlooks details and introduces mistakes.
 Data-Copilot Outperforms Advanced Agent Strategies in Both Accuracy and Efficiency.
-As shown in Table 5, Data-Copilot surpasses all baseline strategies in success rate (Pass@1), and efficiency (#token).
+As shown in Table 3, Data-Copilot surpasses all baseline strategies in success rate (Accuracy), and efficiency (#token).
 Compared to the best baseline (Reflexion), Data-Copilot achieved an +11.2% improvement in accuracy and a -75% reduction in token consumption.
 Besides, as shown in the last two rows of Table 3, Data-Copilot can seamlessly integrate with agent strategies such as ReAct and Reflexion into the workflow deployment.
-For example, when combined with ReAct, Data-Copilot invokes an interface, obtains an intermediate result, and then reasons to invoke the next interface, further improving its performance by +1.3.
+For example, when combined with ReAct, Data-Copilot invokes an interface, obtains an intermediate result, and then reasons to invoke the next interface, improving its performance by +1.3.
 Data-Copilot Reduces Repetitive Generation in the Real World.
-In the real world, numerous requests are similar or even repetitive.
-Data-Copilot can save 70% of the baseline’s consumption because its output is very concise and only contains interface names and arguments.
+In the real world, most requests are similar or even repetitive.
+As shown in Table 3, Data-Copilot can save 70% of token consumption since its output only contains interface names and arguments.
 In contrast, baseline strategies have to repetitively generate complete code for each request, consuming many more tokens.
-Table 6: Different LLM combinations for two stages. Using a weaker LLM for interface design affects both the interface’s quantity and quality.
-Direct Code Generation: GPT-3.5: 28.3, GPT-4: 70
-LLM for Workflow Deployment
+4.3 Why Data-Copilot Brings Improvements?
+Data-Copilot exhibits superior performance in complex scenarios.
+We categorize test requests by entity count: single entity, multiple entities, and multiple entities with complex relations (statistics in Section B.2).
+As Table 4 shows, with single entities, improvement is minimal (4%).
+However, with multiple entities and complex relations, improvements reach 5.3% and 19%, respectively.
+Baselines often generate logically incorrect code and omit critical steps, while Data-Copilot simply invokes versatile interfaces, reducing real-time response complexity through pre-designed interfaces.
 GPT-3.5 GPT-4
-LLM for Interface Design GPT-3.5 31.9 49.6
+Direct Code Generation 28.3 70
+Workflow Deployment LLM
+GPT-3.5 GPT-4
+Interface Design LLM GPT-3.5 31.9 49.6
 GPT-4 70.2 74.6
-Table 7: We ablated 3 designs from Data-Copilot.
-Methods Pass@1(%)
+Table 5: We explore different LLM combinations for interface design and workflow deployment.
+The Quality of Interface is the Key Factor.
+We analyze the effects of using different LLM combinations for two stages: interface design and workflow deployment.
+As shown in Table 5, we observe when using a weaker LLM for interface design, the effectiveness is significantly diminished.
+For example, using GPT-3.5-turbo (1st stage) and GPT-4-turbo (2nd stage) resulted in a score of only 49.6, which is even lower than directly using GPT-4-turbo for direct code generation (70.2).
+This phenomenon can also be seen in other combinations.
+After manually checking, we find interfaces designed by GPT-3.5-turbo are prone to failure and exhibit poor generalizability.
+Therefore, we chose to use GPT-4-turbo for interface design and optimization, while employing GPT-3.5-turbo for real-time deployment, achieving a balance between accuracy and efficiency.
+Methods Accuracy(%)
 Direct Code Generation 28.5
@@
 $w/o$ Interface-Code Hybrid 63.8
-4.3 Analysis: Why Data-Copilot Can Bring Improvements?
-Data-Copilot Exhibits a Pronounced Advantage on More Complex Scenarios.
-To better analyze the behavior of Data-Copilot, we categorize each test request into different subsets based on the number of entities it involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations).
-The statistics for each subset are shown in Section B.1.
-As shown in Table 5, for the single entity subset (denoted as Single), which represents the simplest case, the baselines and our method perform similarly, with only a 4% improvement.
-However, for the more challenging multi-entities (denoted as Multiple) or multi-entities with complex relation (denoted as Complex), our improvements are substantial, with enhancements of 5.3% and 19%, respectively.
-We find that in these complex cases, baseline methods often generate logically incorrect code on the first attempt and may omit some critical steps.
-In contrast, Data-Copilot only needs to invoke versatile interfaces and deploy workflows, significantly reducing the complexity of real-time responses.
-These findings highlight the advantages of Data-Copilot, which reduces task complexity by pre-designing interfaces in advance.
-Interface Quality is the Key Factor.
-We analyze the effects of using different LLM combinations for two stages: interface design and workflow deployment.
-As shown in Table 7, we observed that when using a weaker LLM for interface design, the effectiveness is significantly diminished.
-For example, using GPT-3.5-turbo (1st stage) and GPT-4-turbo (2nd stage) resulted in a score of only 49.6, which is even lower than directly using GPT-4-turbo for code generation (70.2).
-This phenomenon can also be seen in other combinations.
-After manually checking the interfaces in this case, we found these interfaces designed by GPT-3.5-turbo are limited in number and of poor quality.
-Therefore, we chose to use GPT-4-turbo for interface design and optimization, while employing GPT-3.5-turbo for real-time deployment, achieving a balance between accuracy and efficiency.
-4.4 Ablation Study
-As shown in Table 7, we ablate Data-Copilot from the following three aspects:
-  1. 1.
-    Self-exploration Request.
-We ablate the self-exploration request process Section 2.1.
-Instead, we directly used seed requests for interface design and optimization.
-In this case, the performance dropped significantly by 34.3.
-This may be because there are too few seed requests, which are insufficient to design universe interfaces.
-  2. 2.
-    Interface Optimization.
-We ablate the interface optimization phase during interface design.
-It means every successfully designed interface was retained in the interface library.
-We observed that performance was also significantly affected (-28 points).
-Without interface optimization, there are many similar interfaces, which could hurt the effect of workflow invocation.
-  3. 3.
-    Interface-Code Hybrid Generation.
-Data-Copilot only can invoke interfaces during workflow deployment.
-Directly generating code is not allowed even when encountering requests that cannot be resolved by interfaces.
-After ablation, we observe a performance drop of 6.4 points.
-These losses are mainly caused by new requests that cannot be covered by existing interfaces.
-It highlights the flexibility of Data-Copilot, which can flexibly solve familiar requests by invoking interfaces or handle new requests by generating code directly.
+Table 6: We ablate three modules from Data-Copilot.
 5 Conclusion
-We propose a data analysis agent, Data-Copilot, to address a wide array of data-related tasks.
-Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests.
-It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
-Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
-Appendix
-Appendix A Preprocessing of Workflow Deployment
-Intent Analysis
-To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks.
-For example, the request is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2019 to May 2024 for comparison".
-To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
-Multi-form Output
-Upon execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text.
-Additionally, it also provides a comprehensive summary of the entire workflow.
-It greatly enhances the interpretability of the whole process, as the interface workflow is easy for humans to read and inspect.
-As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...".
-Data-Copilot first interprets the user’s intent.
-Then it deploys a three-step workflow: 1) Invoking get-GDP-data() interface to acquire historical GDP data.
-2) Invoking predict-next-value() interface for forecasting.
-3) Visualizing the output.
-Appendix B Evaluation Details
-B.1 The Definitions of Three Complexity Levels for Testset
-We categorize our test set into three types based on the number of entities involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations).
-The ststics of three subset are shown in Table B1.
-  • Single entity: Requests involve a single entity and can be resolved step-by-step. E.g., "query based on a specific condition".
-  • Multiple entities: Requests require processing multiple entities simultaneously. E.g., "compare a certain metric across multiple entities".
-  • Multiple entities with complex relations: Involving multiple entities and containing loops, nesting, and other intricate logic.
-E.g., "List the top 10 stocks by yesterday’s price increase that also in the internet industry."
-Table B1: Statistics of four task types and three complexity levels on our test set.
-Task Types Request Complexity
-Single Entity Multiple Entities Multi-entities with Complex Relation Overall
-Stock 79 106 87 272
-Fund 55 36 22 113
-Corp. 97 30 4 131
-Other 6 12 13 31
-Total 237 184 126 547
-B.2 Human Evaluation on Test Benchmark
-We invite four more graduate students to manually evaluate our benchmark according to four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy.
-The belief guidance for human evaluation is as follows:
-  • Task Difficulty: The difficulty of the task, whether it requires multiple steps… Scoring Criteria: 5: very difficult,…, 1: easy.
-  • Request Rationality: Whether the request is reasonable, or if it is strange and does not align with human habits… Scoring Criteria: 5: Reasonable, aligns with humans..
-  • Expression Ambiguity: Whether the phrasing of the request is ambiguous, or if the entities involved are unclear… Scoring Criteria: 5: Clear without any ambiguity…
-  • Answer Accuracy: Whether the answer table is correct, detailed, and comprehensive, totally meeting user expectations… Scoring Criteria: 5: Data is totally correct … 3: Partial…
-Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
-B.3 Expanding to Other Programming Languages.
-In addition to the Python language, Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces.
-We tested three programming languages: Python, C++, and Matlab.
-The results indicate that Python performed the best (Pass@1: 70.2), followed by C++ (54.2), with Matlab (36.5) yielding the poorest results.
-Upon examination, we found that Matlab code often suffers from formatting errors or dimensional discrepancies in the data, rendering the program non-executable.
-We speculate this may be related to a lack of sufficient Matlab code in the pre-training corpus.
-So ultimately, we opted for Python for Data-Copilot.
-Appendix C Limitations
+We propose Data-Copilot, a code-centric data analysis agent.
+It generates code for large-scale data processing and creates interface modules through data exploration, improving real-time request success.
+It autonomously designs universal interfaces for various data types and invokes them for reliable problem-solving.
+Experiments show higher success rates with lower token consumption.
+Limitations
 Data-Copilot proposes a new paradigm for addressing the data-related task, through LLM. But we want to highlight that it still remains some limitations or improvement spaces:
@@
 Therefore, we believe Data-Copilot also possesses the potential to effectively handle data from other domains.
-Appendix D Related Works
-In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa [Brown et al., 2020, Chowdhery et al., 2022, Zhang et al., 2022, Zeng et al., 2023, Touvron et al., 2023, Ouyang et al., 2022, OpenAI, 2023, Wei et al., 2022a] have revolutionized the field of natural language processing (NLP).
+Appendix
+Appendix A Preprocessing of Workflow Deployment
+Intent Analysis To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks.
+For example, the request is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2019 to May 2024 for comparison".
+To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
+Multi-form Output Upon execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text.
+Additionally, it also provides a comprehensive summary of the entire workflow.
+It greatly enhances the interpretability of the whole process, as the interface workflow is easy for humans to read and inspect.
+As the example shown in Figure F2, the request is "Forecasting China’s GDP growth rate...".
+Data-Copilot first interprets the user’s intent.
+Then it deploys a three-step workflow: 1) Invoking get-GDP-data() interface to acquire historical GDP data.
+2) Invoking predict-next-value() interface for forecasting.
+3) Visualizing the output.
+Appendix B Experiments Details
+B.1 Baselines and Experiments Details
+We compare Data-Copilot with ② direct code generation (Direct-Code) and various agent strategies: ③ ReAct (Yao et al., 2022): LLMs iteratively combine reasoning and code execution.
+④ Reflexion (Shinn et al., 2023): LLMs refine responses based on compiler feedback, limited to two iterations.
+⑤ Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): Three LLM agents (two coders, one manager) collaborate to generate solutions.
+Data-Copilot uses the same LLM as all baselines when invoking pre-designed interfaces.
+Detailed prompts are in Section E.5.
+B.2 The Definitions of Three Complexity Levels for Testset
+We categorize our test set into three types based on the number of entities involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations).
+The ststics of three subset are shown in Table B1.
+  • Single entity: Requests involve a single entity and can be resolved step-by-step. E.g., "query based on a specific condition".
+  • Multiple entities: Requests require processing multiple entities simultaneously. E.g., "compare a certain metric across multiple entities".
+  • Multiple entities with complex relations: Involving multiple entities and containing loops, nesting, and other intricate logic.
+E.g., "List the top 10 stocks by yesterday’s price increase that also in the internet industry."
+Table B1: Statistics of four task types and three complexity levels on our test set.
+Task Types Request Complexity
+Single Entity Multiple Entities Multi-entities with Complex Relation Overall
+Stock 79 106 87 272
+Fund 55 36 22 113
+Corporation 97 30 4 131
+Other 6 12 13 31
+Total 237 184 126 547
+Table B2: Human evaluation on the human-proposed requests and synthesized requests across four dimensions.
+Task Difficulty Request Rationality Expression Ambiguity Answer Accuracy
+Seed Set 3.5 4.3 4.4 -
+Test Set 3.9 4.2 4.0 4.8
+B.3 The Quality of Self-exploration Requests.
+To assess the quality of self-exploration requests, we invited four additional graduate students to manually evaluate our test set (546 requests) and human-proposed requests (173 seed set) on four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy.
+We provide detailed guidance for each criterion in Section B.5.
+The results of two sets are shown in Table B2.
+We observed that the synthesized requests exhibit a comparable quality to human-proposed requests, with slightly higher difficulty, ambiguity, and similar rationality.
+It ensures that our test set can reflect most of the real-world demands.
+Besides, the evaluators gave a high score of 4.8 on the label (answer table) of our test set, which also ensures the accuracy of our dataset.
+B.4 The detail of GPT-4o Evaluation
+Table B3: Comparison of Human and GPT-4o Evaluations
+Human Evaluation GPT-4o Evaluation
+Average Score 65.8 67.2
+Correlation Coefficient 0.894 0.894
+N(score $\geq$ 60) 35 cases 38 cases
+As described in Section 4.1, the evaluation contains three aspects: Data-table, Image, and Efficiency.
+Data-table Evaluation: For each request, we use GPT-4o to compare predicted data table with human-annotated table.
+If GPT-4o identifies any inconsistencies, the judgment is False.
+For example, a request: “Please show me the China’s GDP ….” with its labeled data-table: [2023: 17.8 , 2022: 17.9 , 2021: 17.8 , 2020: 14.6].
+Predicted data: [2023: 17.8 , 2022: 17.9 , 2021: 17.8].
+Based on them, judgment of GPT-4o is False.
+Image Evaluation: We design a comprehensive evaluation checklist for GPT-4o-based image scoring, comprising 5 main categories with 10 sub-dimensions.
+It includes (1) numerical points, (2) lines, (3) axes, (4) aesthetics of image layout, and (5) chart design, e.g., Chart type, Color usage, Proportion and scale, Labels and legends, Readability, Completeness, Relevance, Distinctiveness, Data accuracy, … If the total score exceeds 60 points, the image is considered to meet expectations (True); otherwise, it is judged as false.
+Manual Evaluation Vs.
+GPT-4o Evaluation.
+For data-table evaluation, GPT-4o’s assessment results are highly accurate as it only needs to compare differences in text modality.
+For image evaluation, we randomly sampled 50 examples for both human and GPT-4o-based evaluation.
+As shown in Table B3, we calculate the average score, correlation coefficient, and the #samples >= 60 of the two methods.
+The results show that the average scores of the two evaluation methods are close (GPT-based: 67.2, Human-based: 65.8), and the two score sequences also have a strong correlation (0.894).
+B.5 Human Evaluation on Test Benchmark
+We invite four more graduate students to manually evaluate our benchmark according to four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy.
+The belief guidance for human evaluation is as follows:
+  • Task Difficulty: The difficulty of the task, whether it requires multiple steps… Scoring Criteria: 5: very difficult,…, 1: easy.
+  • Request Rationality: Whether the request is reasonable, or if it is strange and does not align with human habits… Scoring Criteria: 5: Reasonable, aligns with humans..
+  • Expression Ambiguity: Whether the phrasing of the request is ambiguous, or if the entities involved are unclear… Scoring Criteria: 5: Clear without any ambiguity…
+  • Answer Accuracy: Whether the answer table is correct, detailed, and comprehensive, totally meeting user expectations… Scoring Criteria: 5: Data is totally correct … 3: Partial…
+Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
+B.6 Ablation Study
+As shown in Table 6, we ablate Data-Copilot from three aspects: ① We ablate the self-exploration request process.
+Instead, we directly used seed requests for interface design and optimization.
+It leads to a 34.3 performance drop.
+We observe that too few seed requests are insufficient to design universe interfaces.
+② We ablate the interface optimization.
+It means every successfully designed interface was retained in the interface library.
+We observe that performance is also significantly affected (-28 points).
+Without interface optimization, there are many similar interfaces, which hurt the effect of workflow invocation.
+③ Interface-Code Hybrid Generation: Data-Copilot can only invoke interfaces during workflow deployment.
+Without a hybrid generation manner, it shows a 6.4-point performance drop, which is caused by “uncovered” requests.
+It highlights our flexibility in addressing different types of requests.
+B.7 Expanding to Other Programming Languages
+In addition to the Python language, Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces.
+We tested three programming languages: Python, C++, and Matlab.
+The results indicate that Python performed the best (Accuracy: 70.2), followed by C++ (54.2), with Matlab (36.5) yielding the poorest results.
+Upon examination, we found that Matlab code often suffers from formatting errors or dimensional discrepancies in the data, rendering the program non-executable.
+We speculate this may be related to a lack of sufficient Matlab code in the pre-training corpus.
+So ultimately, we opted for Python for Data-Copilot.
+Appendix C Related Works
+In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023; Ouyang et al., 2022; OpenAI, 2023; Wei et al., 2022a; Zhang et al., 2024a) have revolutionized the field of natural language processing (NLP).
 These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning.
-The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism [Ouyang et al., 2022, Wang et al., 2022b, a].
+The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism (Ouyang et al., 2022; Wang et al., 2022b, a).
 LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs).
-Tool-enhanced studies  [Schick et al., 2023, Gao et al., 2022, Qin et al., 2023a, Hao et al., 2023, Qin et al., 2023b] integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools.
-Several researchers have extended the scope of LLMs to include the other modality [Wu et al., 2023a, Surís et al., 2023, Shen et al., 2023, Liang et al., 2023b, Huang et al., 2023].
-In addition, there are many LLM-based agent applications [Xie et al., 2023], such as CAMEL [Li et al., 2023a], AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^3] https://github.com/reworkd/AgentGPT, BabyAGI [^4] https://github.com/yoheinakajima/babyagi, BMTools [^5] https://github.com/OpenBMB/BMTools, LangChain [^6] https://github.com/hwchase17/langchain, Agentverse [Chen et al., 2023b], Autoagent [Chen et al., 2023a], MetaGPT [Hong et al., 2023], AutoGEN [Wu et al., 2023b], etc.
+Tool-enhanced studies (Schick et al., 2023; Gao et al., 2022; Qin et al., 2023a; Hao et al., 2023; Qin et al., 2023b; Hou et al., 2023) integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools.
+Several researchers have extended the scope of LLMs to include the other modality (Wu et al., 2023a; Surís et al., 2023; Shen et al., 2023; Liang et al., 2023b; Huang et al., 2023).
+In addition, there are many LLM-based agent applications (Xie et al., 2023), such as CAMEL (Li et al., 2023a), AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^3] https://github.com/reworkd/AgentGPT, BabyAGI [^4] https://github.com/yoheinakajima/babyagi, BMTools [^5] https://github.com/OpenBMB/BMTools, LangChain [^6] https://github.com/hwchase17/langchain, Agentverse (Chen et al., 2023b), Autoagent (Chen et al., 2023a), MetaGPT (Hong et al., 2023), AutoGEN (Wu et al., 2023b), etc.
 Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks.
-Except for learning to operate the tools, several contemporaneous studies [Cai et al., 2023, Qian et al., 2023] have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning.
+Except for learning to operate the tools, several contemporaneous studies (Cai et al., 2023; Qian et al., 2023) have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning.
 These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
-Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers.
-FLAME [Joshi et al., 2023] investigates the feasibility of using NLP methods to manipulate Excel sheets.
-StructGPT [Jiang et al., 2023] explore reasoning abilities of LLM over structured data.
-LiDA [Dibia, 2023] and GPT4-Analyst [Cheng et al., 2023] focus on automated data exploration.
-Besides, many reseraches [Liu et al., 2023, Chang and Fosler-Lussier, 2023, Dong et al., 2023], like Sheet-Copilot [Li et al., 2023b], BIRD [Li et al., 2023c], DAIL-SQL [Gao et al., 2023], DIN-SQL [Pourreza and Rafiei, 2023], PET-SQL [Li et al., 2024b], DB-Copilot [Wang et al., 2023b], MAC-SQL [Wang et al., 2023a], ACT-SQL [Zhang et al., 2023], ChatBI [Lian et al., 2024], CodeS [Li et al., 2024a], SQLPrompt [Sun et al., 2023a], ChatDB [Hu et al., 2023], SQL-PaLM [Sun et al., 2023b], and DB-GPT [Xue et al., 2023], EHRAgent [Shi et al., 2024] apply LLMs to Text2SQL and table rasoning.
-Chain-of-Table [Wang et al., 2024] proposes a step-by-step reasoning strategy based on the table.
-Distinct from these approaches, our Data-Copilot system is different in the following aspects:
-(1) Different Task Settings: Unlike Text2SQL methods, Data-Copilot focuses more on how to utilize data and present it according to user preferences, rather than retrieving data from databases.
-We assume that Data-Copilot has access to a vast amount of data, which is available via API.
-(2) Richer Usage Scenarios: Data-Copilot needs to meet a diverse range of human needs, such as forecasting and plotting.
-Therefore, it employs programmable language like Python for data processing, offering more varied use cases than SQL.
-(3) Distinct Implementation Approach: Data-Copilot is a code-centric agent for data processing, rather than relying on SQL or pre-existing tools.
-This approach is more flexible and can meet complex user requirements.
-Appendix E Visualization
+Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers (Maddigan and Susnjak, 2023; Valverde-Rebaza et al., 2024; Gu et al., 2024a; Liu et al., 2024b; Chen et al., 2024a; Zhang et al., 2023b; Ahn, 2024; Inala et al., 2024; Liu et al., 2024a; Xie et al., 2024; Wu et al., 2024; Guo et al., 2024a; Cao et al., 2024; Lu et al., 2024; Ye et al., 2024; Sui et al., 2024; Ford et al., 2024; Chen et al., 2024b; Weng et al., 2024; Shen et al., 2024).
+FLAME (Joshi et al., 2023) investigates the feasibility of using NLP methods to manipulate Excel sheets.
+StructGPT (Jiang et al., 2023) explore reasoning abilities of LLM over structured data.
+LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023; Ma et al., 2023) focus on automated data exploration.
+Besides, many reseraches (Liu et al., 2023; Chang and Fosler-Lussier, 2023; Dong et al., 2023; Almheiri et al., 2024), like Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2024b), DAIL-SQL (Gao et al., 2023), DIN-SQL (Pourreza and Rafiei, 2023), PET-SQL (Li et al., 2024c), DB-Copilot (Wang et al., 2023b), MAC-SQL (Wang et al., 2023a), ACT-SQL (Zhang et al., 2023a), ChatBI (Lian et al., 2024), CodeS (Li et al., 2024a), SQLPrompt (Sun et al., 2023a), ChatDB (Hu et al., 2023), SQL-PaLM (Sun et al., 2023b), and DB-GPT (Xue et al., 2023), EHRAgent (Shi et al., 2024) apply LLMs to Text2SQL and table rasoning.
+Chain-of-Table (Wang et al., 2024) proposes a step-by-step reasoning strategy based on the table.
+Some researchers also focus on designing various benchmarks and evaluation methods (Lai et al., 2023; Zhang et al., 2024b; Sahu et al., 2024; Yang et al., 2024; Ma et al., 2024; Gu et al., 2024b; Jing et al., 2024; Hu et al., 2024) for LLMs in data science.
+Appendix D Visualization
 We provide several cases in this section to visualize workflow deployed by Data-Copilot, which includes queries about diverse sources (stocks, company finance, funds, etc.) using different structures (parallel, serial, and loop Structure).
-Different Structures
-As shown in Figure G4,G5, Data-Copilot deploys different structural workflows based on user requirements.
-In Figure G4, the user proposes a complex request, and Data-Copilot deploys a loop structure to implement the financial data query of each stock, and finally outputs a graph and table in parallel.
-In Figure G5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas.
+Different Structures As shown in Figure F4,F5, Data-Copilot deploys different structural workflows based on user requirements.
+In Figure F4, the user proposes a complex request, and Data-Copilot deploys a loop structure to implement the financial data query of each stock, and finally outputs a graph and table in parallel.
+In Figure F5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas.
 These concise workflows can cope with such complex requests well, which suggests that the data exploration and workflow deployment process of Data-Copilot are rational and effective.
-Diverse Sources
-Figure G6, G7, G8 demonstrate that Data-Copilot is capable of handling a large number of data sources, including stocks, funds, news, financial data, etc.
+Diverse Sources Figure F6, F7, F8 demonstrate that Data-Copilot is capable of handling a large number of data sources, including stocks, funds, news, financial data, etc.
 Although the formats and access methods of these data types are quite different, our system efficiently manages and displays the data through its self-designed versatile interface, requiring minimal human intervention.
-Appendix F Detailed Prompts
+Appendix E Detailed Prompts
 We provide detailed prompts for our Data-Copilot and baselines.
-Specifically, for the data exploration, we provide detailed prompts in Section F.1 for four phases: self-exploration, interface design, and interface optimization.
-For the workflow deployment phase, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow.
-Additionally, we outline prompts for all baselines in Section F.3: Direct-Code, ReAct, Reflection, and Multi-Agent Collaboration strategies.
-F.1 Prompts For Data Exploration
+Specifically, for the data exploration, we provide detailed prompts in Section E.3 for four phases: self-exploration, interface design, and interface optimization.
+For the workflow deployment phase, we also provide prompts for three key procedures in Section E.4: Intent Analysis, Task Selection, and Planning Workflow.
+Additionally, we outline prompts for all baselines in Section E.5: Direct-Code, ReAct, Reflection, and Multi-Agent Collaboration strategies.
+E.1 Algorithm Flow For Interface Design and Optimization
+Step 1: Interface Design and Testing
+1. One-to-One Interface Design
+* Design interfaces one-by-one using synthesized requests (request 1----> interface 1)
+* Generate a series of initial interfaces with single functionalities
+2. Test Case Generation
+* For each designed interface:
+> Sample a sub data-table from relevant data sources
+> Generate test requests based on data schemas and sampled sub-tables
+> For example: For an interface: "GDP_retrieves(Year, Country)"
+- Sample a data point from the GDP table: China, 2023, \$13 trillion
+- Generate test request: "What is Chinas GDP in 2023?"
+- Use the sampled sub-table as the expected answer
+> Repeat this process K times for each interface, generating K test cases
+3. Interface Evaluation
+* Use the generated test cases to evaluate the usability of each interface
+* Record interfaces that pass the tests and their testing results
+Step 2: Interface Optimization and Compiler Feedback Evaluation
+1. Similar Interface Retrieval
+* Analyze the functionality and parameters of all interfaces
+* Identify pairs of interfaces with similar or overlapping functionalities
+2. Interface Merging Decision
+* For each pair of similar interfaces, assess the necessity and feasibility of merging
+* Consider functional coverage, usage scenarios, and complexity of the interfaces
+3. Interface Merging Execution
+* Design a new merged interface ensuring:
+> The new interface covers all functionalities of the original two interfaces
+4. Merged Interface Evaluation
+* Evaluate the merged interface using test cases generated in the previous steps
+* Collect compiler feedback on the merged interface
+* Verify:
+> Whether the new interface can correctly handle all test cases of the original interfaces
+> Whether the output is consistent with the original interfaces in both format and content
+5. Feedback-Based Self-Optimization
+* If the merged interface fails evaluation:
+> Guide the LLMs to analyze compiler feedback
+> Perform self-reflection based on the feedback and two old interfaces
+> Optimize the interface design to address issues
+* Repeat the evaluation-reflection-optimization cycle multiple times
+* Abandon the merging attempt after multiple fails
+E.2 Example for Data Exploration
+###Instruction: Given some data and its description, please mimic these seed requests and generate more requests.
+The requests you generate should be as diverse as possible, covering more data types and common needs.
+###Seed Request: {request1, request2,...}
+###Parsing file for GDP_Data:{
+Description: This data records China’s annual and quarterly GDP...,
+Access Method: pro.cn-gdp(start-time, end-time, frequency,...),
+Output Schema: Return 9 columns, including quarter, gdp, gdp-yoy...,
+Usage:{
+Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’,..),
+First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
+###Parsing file for Stock_data:{...}, ....
+E.3 Prompts For Data Exploration
 Explore Data by self-exploration Phase
@@
 Interfaces after merging: {interface1, interface2, optimized interface3, ..}
-F.2 Prompts For Workflow Deployment
+E.4 Prompts For Workflow Deployment
 Intent Analysis Phase
@@
 Please output an interface invocation for this instruction.
-F.3 Prompts For Baselines
+E.5 Prompts For Baselines
 Direct-Code LLM
@@
 ###Manager Decision: {Final Solution}
-Appendix G Case Study
-Detailed cases of data exploration and workflow deployment are presented in Sections G.1 and G.2.
-G.1 Cases for Data Exploration
+Appendix F Case Study
+Detailed cases of data exploration and workflow deployment are presented in Sections F.1 and F.2.
+F.1 Cases for Data Exploration
 ---- The First Request: Interface Design -----
@@
 {Interface3=Interface Name:plotData(), ...}
-G.2 Cases for Real-time Request
+F.2 Cases for Real-time Request
 ### Request: Today is 2019-03-13. The user request is to please help me plot the stock price trend of Guizhou Maotai from January 23, 2018 to today.
@@
 "description": "Guizhou Maotai Stock Price Line Chart"
-Figure G2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
-Figure G3: We have listed some of Data-Copilot’s self-design interfaces, including five categories.
+Figure F2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
+Figure F3: We have listed some of Data-Copilot’s self-design interfaces, including five categories.
 Most of the interfaces are a combination of several simple interfaces.
 For example, the text marked in blue indicates that the interface includes multiple functions.
-Figure G4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
-Figure G5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
-Figure G6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
-Figure G7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
-Figure G8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
-Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.
+Figure F4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
+Figure F5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
+Figure F6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
+Figure F7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
+Figure F8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
+Figure F9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.